### v2

# Talent Hoarding in Organizations

Abstract
Most organizations rely on managers to identify talented workers. However, managers who are evaluated on team performance have an incentive to hoard workers. This study provides the first empirical evidence of talent hoarding using personnel records and survey evidence from a large manufacturing firm. Talent hoarding is reported by three-fourths of managers, is detectable in managerial decisions, and occurs more frequently when hoarding incentives are stronger. Using quasi-random variation in exposure to talent hoarding, I demonstrate that hoarding deters workers from applying to new positions, inhibiting worker career progression and altering the allocation of talent in the firm.
1 Introduction
Firms must continually decide how to allocate workers to jobs, a process which has critical implications for productivity (Rosen, 1982, Holmstrom and Tirole, 1989). Because it is difficult to perfectly observe worker ability, most firms rely on managers to identify talented workers who can be promoted to higher-level positions. However, when a talented worker leaves their team for a new position, team performance suffers. Since managers are rewarded based on team performance and firms cannot perfectly monitor manager actions, the conflicting interests of manager and firm create the potential for moral hazard (Holmstrom, 1979). A growing body of evidence documents that workers in high-level positions have large impacts on firm performance (Bloom and Van Reenen, 2007, Lazear et al., 2015), implying that managers may create significant efficiency costs if they hoard talented workers rather than promote them. Ample anecdotal and survey evidence points to widespread talent hoarding in organizations. In a global survey, half of organizations report that managers hoard talent by discouraging worker mobility (i4cp, 2016). A US-based survey finds that workers in one-third of firms feel the need to keep internal applications secret from their managers out of fear of retaliation (KornFerry, 2015). In Germany, 83% of the top publicly listed companies cite managerial talent hoarding as a key friction in their organization (hkp, 2021). [^2] Talent hoarding also occurs in science and academia. Zuckerman (2021) documents how Katalin Karikó, a seminal developer of mRNA vaccines and 2023 Nobel prize winner, experienced talent hoarding when she decided to leave her lab for a new position. Her advisor Robert Suhadolnik vowed “to do whatever he could to stop his protégée from leaving…he made it clear she had two career choices. ‘You can work in my lab or go home,’ he told her. Suhadolnik followed through on this threat, telling a local immigration office that she was living in the country illegally and should be deported.” Despite the apparent prevalence of talent hoarding and its likely detrimental consequences, very little empirical evidence on talent hoarding exists in economics. Studying talent hoarding empirically is challenging. Managers often hoard talent through hidden actions that are difficult to observe, even in rich datasets. Furthermore, identifying the causal impacts of talent hoarding requires plausibly exogenous variation in hoarding. This study provides the first empirical evidence on talent hoarding in organizations. I combine rich personnel records from a large manufacturing firm with survey evidence from both managers and workers in the firm. I document widespread managerial incentives to hoard talent: the majority of managers perceive a conflict of interest with respect to developing talent, because more developed workers are more likely to leave the team. Talent hoarding is reported by 75% of managers. Using a direct measure of talent hoarding derived from the firm’s personnel records, I test predictions from my conceptual framework and confirm that stronger hoarding incentives lead to more talent hoarding. I leverage quasi-random variation in workers’ exposure to talent hoarding to demonstrate the impacts of talent hoarding on workers’ career outcomes. Hoarding significantly deters workers’ applications for internal positions, many of which would have been successful. Because talent hoarding has a larger impact on high-quality workers, hoarding reduces the quality of the applicant pool, thus restricting the firm’s ability to effectively allocate talent within the organization.
I develop a simple conceptual framework that formalizes the misaligned incentives that give rise to talent hoarding. Since managers are both tasked with identifying productive workers and are rewarded based on team productivity, managers have an incentive to hoard talent by preventing workers from seeking new positions outside the team. This framework yields empirical predictions for both manager and worker behavior. For example, when gains to hoarding are larger or worker departure from the team is costlier, managers engage in more hoarding. High-productivity workers who experience more hoarding are more deterred from applying, which negatively impacts internal applicant pools and has the potential to distort the allocation of talent within the firm. To empirically test for talent hoarding, I collect a unique combination of personnel records, internal job application data, and new survey evidence from a large manufacturing firm that employs over 200,000 workers. I focus my analysis on the firm’s largest internal labor market, consisting of over 30,000 white-collar and management employees in Germany. I demonstrate that the firm is similar to other large firms in terms of its workforce and organizational design. As in many other large firms, employees who want to switch to a new position or to be promoted are required to apply for the position. Since most workers in the same team work in similar positions, workers who seek different positions, including promotions, typically need to apply to jobs outside the team. The data I assemble offer three key advantages. First, collecting new survey evidence from managers in the firm allows me to directly document that managers face misaligned incentives and act accordingly. Second, the firm’s personnel records allow me to construct a measure of talent hoarding using the extent to which managers systematically suppress the visibility of workers. Without such data, directly measuring managerial talent hoarding is empirically challenging because hoarding by definition involves hidden actions. Third, by combining personnel records with the universe of application and hiring decisions at the firm, I am able to document the extent to which talent hoarding deters workers’ high-stakes applications that would otherwise have been successful. Survey responses from 62% of managers in the firm reveal that incentives to hoard talent are widespread. The majority of managers (55%) explicitly acknowledge that supporting workers’ career development entails a conflict of interest because more developed workers are more likely to leave the team. This conflict of interest relates to the misaligned incentives managers face: while almost all managers (96%) believe they would have a large impact on workers’ careers, only 36% perceive this impact to be valued by the firm. A further disincentive to develop talent is the risk of workers leaving the team: 68% of managers state that they would be more likely to support workers’ career development if it would be easier to replace lost talent. The survey demonstrates that managerial talent hoarding actions, which are reported by 75% of managers, are a common response to these prevailing managerial incentives. Managers’ revealed actions provide further evidence that managers purposefully hoard talent. To capture talent hoarding in the firm data, I construct a direct hoarding measure based on managers’ systematic suppression of worker visibility using two worker ratings that managers conduct. The ratings differ in the extent of their visibility outside of a worker’s team, meaning that managers can suppress worker visibility by lowering the public rating relative to the private rating. Such suppression of visibility is cited as a key form of talent hoarding in both managers’ and workers’ survey responses. Testing predictions from my conceptual framework based on three proxy measures for managerial hoarding incentives, I demonstrate that managers are indeed more likely to engage in talent hoarding if they face greater incentives to hoard. Additional evidence from alternative measures of talent hoarding based on managers restricting access to high-visibility trainings and managers’ explicit hoarding statements from the survey further corroborates this finding. I use quasi-random variation in workers’ exposure to talent hoarding based on manager rotations to demonstrate that managerial talent hoarding has meaningful impacts on worker outcomes. When a manager learns that they will move to a new position on a different team, they no longer have an incentive to hoard workers on their current team. For workers whose manager will soon rotate, this creates a temporary window of time during which they are not subject to talent hoarding. Consistent with the cessation of talent hoarding, I find that manager rotations increase worker applications by 78%. To support the interpretation of rotation effects as being driven by talent hoarding, I test predictions from the conceptual framework, which identifies specific subgroups of workers who should face higher levels of hoarding and therefore exhibit larger responses to manager rotations. Coupled with a series of tests for potentially confounding mechanisms, these empirical exercises support my interpretation based on talent hoarding and motivate the use of manager rotations to measure the impacts of talent hoarding on worker outcomes. Talent hoarding has important consequences for workers’ careers and the firm’s ability to allocate talent. To quantify how successful deterred applications would have been, I instrument for workers’ applications with manager rotations. Marginal applicants (identified via standard IV assumptions), who would not have applied in absence of a manager rotation and only apply because their manager rotates, face a 49.1% likelihood of landing a new position. This large and positive hiring likelihood suggests that many deterred applicants would have been successful and is particularly striking given that the average hiring likelihood is 27.6%. The positive selection of marginal applicants is further corroborated by a complier analysis (Abadie, 2003), which implies that talent hoarding has negative impacts on the composition of the applicant pool. Using workers’ performance ratings, I present evidence suggesting that a substantial share of deterred applicants would have been highly productive in their new positions. Together, my findings suggest that in the presence of talent hoarding, firms may forgo productivity gains from hoarded workers who are not assigned to positions in which they would have been productive. Talent hoarding may also exacerbate between-group inequality at the firm. A survey of the firm’s employees suggests that male and female workers may react differently to talent hoarding. In line with the literature on gender differences in preferences (Bertrand, 2011), the survey finds that women in the firm are 22% more likely to place a high value on preserving a good relationship with their manager and are 26% more likely to rely on managers’ career guidance when making career decisions. The effects of manager rotations corroborate that talent hoarding has differential effects by gender, affecting women at a higher part of the quality distribution compared to men. In particular, marginal female applicants are more qualified in terms of their educational qualifications and past performance. They are also more positively selected with respect to their hiring probability for higher-level positions. When comparing potential outcomes for marginal applicants, I find that both men and women experience higher earnings if they choose to apply in the absence of talent hoarding; however, the larger gains realized by women lead to a reduction in the pay gap of 86%. A number of factors suggest that talent hoarding is very likely to manifest similarly in other organizations. The firm I study is similar to other large firms in Germany both in terms of the characteristics of its workforce and its internal organization, in that it is standard that managers are tasked with identifying talented workers, but are neither monitored nor rewarded in doing so (hkp, 2021). Companies across the world report that talent hoarding is commonplace, creates barriers to talent allocation, and occurs through many of the same managerial behaviors that are documented in this study (i4cp, 2016, KornFerry, 2015, Matuson, 2015, Sullivan, 2017). This study contributes to two broad strands of research on organizations. Prior theoretical research has hypothesized that managers engage in self-interested behavior (Holmstrom and Tirole, 1989), largely focusing on managers’ misaligned incentives in the context of biased performance evaluations (Milgrom and Roberts, 1988, Prendergast and Topel, 1996, Fairburn and Malcomson, 2001). While research studying internal labor markets has documented the importance of incentive provision in organizations (Gibbons and Waldman, 1999, Prendergast, 1999), little attention has been paid to how managers’ incentive problems may affect the efficiency of job assignments. One notable exception is theoretical work by Friebel and Raith (2022) who show that different organizational designs may change managers’ incentives to train subordinates and accurately represent their abilities. My study provides the first empirical demonstration of a costly moral hazard problem among managers that likely affects the efficient allocation of talent within organizations. Second, a large empirical literature in economics studies the impacts of managers on firm outcomes. The majority of this literature has focused on upper management, and in particular on documenting the impacts of CEOs on firm performance (Bertrand and Schoar, 2003, Bennedsen et al., 2007, Malmendier and Tate, 2009, Bennedsen et al., 2020). An emerging body of work uses detailed data on managers and workers to show that even managers at lower levels of the firm hierarchy have large impacts on worker outcomes, including worker productivity (Lazear et al., 2015, Frederiksen et al., 2020, Fenizia, 2022), turnover (Hoffman and Tadelis, 2021), and career progression (Kunze and Miller, 2017, Cullen and Perez-Truglia, 2023, Benson et al., 2021). This study adds to these findings by uncovering talent hoarding as an important mechanism that influences managers’ impacts on firms and workers. By demonstrating that talent hoarding has meaningful impacts on career progression, this study also contributes to both theoretical and empirical work seeking to understand the dynamics of internal labor markets (Waldman, 1984, Milgrom and Oster, 1987, Baker et al., 1994, Benson et al., 2019, Huitfeldt et al., 2023). The rest of the paper proceeds as follows. Section 2 provides anecdotal evidence and introduces a simple framework that offers a formal definition of talent hoarding. Section 3 describes the institutional setting and data. Section 4 focuses on managers and provides direct evidence on the hoarding incentives they face and the actions that follow. Section 5 uses quasi-random variation centered around manager rotations to document the impact of managerial talent hoarding on workers. Section 6 presents evidence on the consequences of talent hoarding for workers’ careers and the firm’s ability to allocate talent. Section 7 concludes.
2 Background and Conceptual Framework
This section presents anecdotal evidence of how talent hoarding occurs in practice. I conduct a large-scale survey at the firm I study, which illustrates the reasons why and the actions through which managers hoard talent. Building on this intuition, a simple conceptual framework offers a formal definition of talent hoarding as well as a set of predictions that guides my empirical analysis. What is talent hoarding? A 2017 Wall Street Journal article describes talent hoarding as “manager’s natural tendency to hold on to top performers instead of working to promote them or transfer them to other areas of the company." (Lublin, 2017). Media outlets also describe how managers hoard talent. A 2015 Forbes article observes that managers who hoard talent “never recommend…people for a promotion in another department." (Matuson, 2015). The industry publication Talent Management & HR writes in 2017 that “hoarding managers, in order to reduce the internal visibility of their top team members, may purposely restrict them from serving on task forces and outside-of-function committees." (Sullivan, 2017). Other hoarding strategies are described as underrating potential for higher-level positions or threatening workers who try to leave the team. Talent hoarding is widespread and occurs in a variety of settings. In a survey of 665 global organizations, covering both the private and public sector, half of organizations report that their managers hoard talent by discouraging worker mobility (i4cp, 2016). In Germany, 83% of the top publicly listed companies cite talent hoarding as a crucial friction in their organization (hkp, 2021). Managers’ talent hoarding actions seem consequential. In one-third of US firms, workers feel the need to keep applications secret from their managers out of fear of retaliation (KornFerry, 2015).
2.1 Evidence from an Employee Survey within the Firm
To provide the first detailed evidence on the dynamics of talent hoarding, I conduct a large-scale survey at the firm I study. All employees in my sample were invited via e-mail by the firm’s HR department and were asked to provide their perspectives on the internal labor market at the firm. Employees described challenges regarding their internal career progression both in the form of multiple-choice answers and in free-text responses. [^3] For questions requiring free-text responses, I exclude respondents from the analysis who did not provide any answer. An abbreviated version of the survey instrument is presented in Appendix Section D.2. The survey received over 15,000 responses, yielding a 50% response rate. Respondents are similar to the overall population in terms of their demographics (Appendix Table D1). [^4] Because of strict data protection regulations in Germany, I am not able to link the survey responses to the administrative employee records at the individual level. This limitation precludes a direct comparison of respondents to non-respondents in the administrative data. Respondents report a variety of different actions through which managers hoard talent, which include suppressing public signals of worker ability, restricting access to high-visibility projects or trainings, and explicitly discouraging workers from applying to internal positions. Appendix Table B1 provides selected quotations. In addition, when asked to state the biggest challenge to their internal career progression, the modal answer (provided by 22% of workers who respond to this question) is managers’ limited support for career progression, such as refusal to assist in career planning and restricted access to trainings that would increase workers’ visibility outside the team. Managers’ actions appear to strongly influence workers’ application decisions. 41% of respondents indicate that they are afraid to apply to internal positions, fearing negative repercussions if managers find out about the application. 25% of workers state that they feel the need to ask for managers’ permission before applying for an internal job opening, even though the firm’s internal policies are meant to enable workers to initiate an application on their own. 16% of respondents indicate that applying away from the team is seen as disloyal. These findings suggest that fear of retaliation represents a key dimension through which talent hoarding deters worker applications.
2.2 A Simple Framework of Talent Hoarding
To formally define talent hoarding, consider a firm that employs two types of agents, managers $m$ and workers $i$. For simplicity, teams are composed of one manager and one worker. Workers are characterized by a marginal productivity $\alpha_{i}$ drawn from some known distribution $G$. The firm seeks to efficiently allocate talent to maximize total firm productivity by choosing which workers to promote to managerial positions. Consistent with Rosen (1982), productivity is maximized when high-ability workers are assigned to high-level positions. Thus, in the absence of any constraints, the firm would fill a new managerial vacancy with the most productive worker and would fill the worker’s vacated position by hiring a worker from outside of the firm (i.e., a random draw from $G$). In practice, firms neither perfectly observe worker productivity, nor do they know which workers would accept a promotion. Accordingly, managers are tasked with identifying high-productivity workers and encouraging them to seek promotions. However, if a high-productivity worker is promoted and leaves their team, that team incurs a productivity loss. Managers, who observe the productivity of their workers, are compensated according to team productivity, creating a conflict of interest between the firm and managers. Talent hoarding is defined as the actions taken by managers that lower the likelihood that a worker applies for and receives a promotion. [^5] Respondents to the survey discussed in Section 2.1 report that managers can deter workers from applying by explicitly discouraging or threatening them, underrating worker ability, and restricting access to high-visibility projects or training. In theory, the firm could offer managers a promotion-contingent contract to resolve the misaligned incentives. In practice, firms generally do not compensate managers for promoting their workers, plausibly because of the practical challenges associated with these contracts (discussed in detail in Friebel and Raith, 2022). The one-period framework proceeds as follows. A managerial vacancy opens exogenously. $M$ managers observe the productivity of the worker on their teams and decide to what extent to engage in talent hoarding. Based on managers’ choices, workers decide whether to apply for a promotion. The firm observes noisy signals of worker productivity (e.g., by conducting interviews) and chooses the worker with the highest signal to fill the vacancy. The promoted worker’s previous position is replaced with an outside hire. Team productivity is realized and managers are compensated. Let $\beta\in[0,\infty)$ index the degree of talent hoarding chosen by a manager, with 0 representing no talent hoarding. [^6] Survey responses presented in Appendix Table B1 indicate that suppression of potential ratings and pressure to refrain from applying are common examples of talent hoarding that can be represented by $\beta$. In Section 4.2, I construct a direct measure of talent hoarding by comparing the private performance ratings to the public potential ratings that managers assign to workers, where $\beta$ can be interpreted as reflecting the disparity between these signals. Let $q(\alpha_{i},\beta)$ denote the equilibrium probability that a worker with productivity $\alpha_{i}$ gets promoted, conditional on applying for a promotion. This promotion probability increases with worker productivity ($\frac{\partial q}{\partial\alpha_{i}}>0$), but decreases in the level of talent hoarding ($\frac{\partial q}{\partial\beta}<0$), and reflects the noisy signal received by the firm. Thus, one can interpret talent hoarding as impacting workers through the likelihood that they get promoted. [^7] In practice, workers report that managers diminish their visibility, thus lowering their promotion prospects. In theory, talent hoarding can also operate through the cost of applying, which would yield similar predictions. Furthermore, I assume that the effect of talent hoarding on the conditional promotion probability is larger for more productive workers ($\frac{\partial^{2}q}{\partial\beta\partial\alpha_{i}}<0$). This assumption relates to the noisy signals of applicant productivity observed by the firm. Intuitively, since low-productivity workers are less likely than high-productivity workers to generate a favorable signal, there is less scope for talent hoarding to lower the likelihood that a low-productivity worker gets chosen among multiple applicants. One example in which this would be the case is if a firm employs a two-part screening strategy in which it chooses a subset of applicants to interview based on the CVs of all applicants. A very low-productivity worker may never clear the bar to be interviewed. Thus, the manager can do little to further lower their hiring likelihood. Workers decide whether to apply by weighing expected costs and benefits. Let $b$ denote the benefits of a successful application and $c$ denote the costs of applying. Workers apply if
$q(\alpha_{i},\beta)b$ $\geq c+\varepsilon_{i}$ (1)
where $\varepsilon_{i}\sim\Psi$ captures worker-specific heterogeneity, with $\Psi$ known to the manager. Therefore, from the manager’s perspective, the probability that the worker leaves the team is given by
$p(\alpha_{i},\beta)=q(\alpha_{i},\beta)\Psi\big{(}q(\alpha_{i},\beta)b-c\big{)}$
[…]
$\max_{\beta}\hskip 2.84526pt(1-p(\alpha_{i},\beta))\alpha_{i}+p(\alpha_{i},\beta)\bar{\alpha}-\frac{\phi_{m}}{2}\big{(}p(\alpha_{i},0)-p(\alpha_{i},\beta)\big{)}^{2}$
This optimization problem yields the following first-order condition, which provides a formal definition of talent hoarding as well as predictions with respect to the realized level of talent hoarding:
$p(\alpha_{i},0)-p(\alpha_{i},\beta^{*})=\frac{1}{\phi_{m}}(\alpha_{i}-\bar{\alpha})$
Definition. (Talent hoarding) When worker $i$’s productivity exceeds the expected productivity of an outside hire, the manager optimally hoards talent by choosing $\beta^{*}>0$. As a result, the likelihood that a worker leaves the team is lower relative to $\beta=0$.
This definition yields following predictions for managers’ talent hoarding actions: Prediction 1. (Worker heterogeneity)
High-productivity workers (i.e., high $\alpha_{i}$) experience more talent hoarding. [^10] In many related models (e.g., Garicano, 2000), ability is multi-dimensional. To the extent that productivity on the worker’s current team $\alpha_{i}$ is positively related with performance external to the current team (e.g., leading a team), managers will have an incentive to hoard talent that conflicts with the firm’s objectives. Prediction 2. (Team heterogeneity)
Workers that are more difficult to replace (i.e., if $\bar{\alpha}$ is lower) experience more talent hoarding. Prediction 3. (Manager heterogeneity)
Talent hoarding is greater among managers with low utility costs of hoarding (i.e., low $\phi_{m}$).
In addition, workers’ decision rule implies that talent hoarding reduces the number of applicants and the quality of the applicant pool, limiting the firm’s ability to promote a high-productivity worker and generating misallocation of talent. Prediction 4. (Number of applicants)
[…]
Prediction 5. (Composition of applicants)
Talent hoarding has larger impacts for high-productivity workers. Therefore, higher levels of talent hoarding lead to a lower-quality applicant pool: If $\alpha_{1}<\alpha_{2}$ and $\beta_{1}<\beta_{2}$implies $\frac{\text{Pr}[i\text{ applies}|\alpha_{2},\beta_{1}]}{\text{Pr}[i\text{ applies}|\alpha_{1},\beta_{1}]}>\frac{\text{Pr}[i\text{ applies}|\alpha_{2},\beta_{2}]}{\text{Pr}[i\text{ applies}|\alpha_{1},\beta_{2}]}$
Appendix Section C contains formal derivations of these predictions.
2.3 Implications for Empirical Tests
The predictions that follow from the conceptual framework suggest that the existence of talent hoarding can be inferred empirically from both manager and worker actions. First, the framework predicts that managers should engage more in talent hoarding if they face stronger incentives to hoard, for instance because of larger financial gains from hoarding. This prediction motivates empirical tests that relate managers’ revealed hoarding actions to variation in their hoarding incentives. This approach has the advantage that, relative to analyzing worker behavior, it is less likely to be confounded by potential worker inaction that may be driven by factors unrelated to talent hoarding. However, this test requires a direct measure of managerial talent hoarding actions, which is empirically challenging to construct given that talent hoarding involves hidden actions by definition. Moreover, since managerial hoarding can take many forms, any one measure can only capture a subset of talent hoarding actions. Second, the framework predicts that workers should be deterred from applying to positions outside the team if they are subject to talent hoarding. Accordingly, an empirical test based on workers requires both observing revealed worker actions and variation in workers’ exposure to managerial hoarding. Using high-stakes worker actions, such as internal applications or job transitions, [^11] Although the language used in the framework is centered around applications for promotions, the intuition and predictions also pertain to any type of internal application to a position outside of the team. From the manager’s perspective, both promotions and lateral moves have the same impacts on team productivity. From the firm’s perspective, deterring both types of transitions can create efficiency costs. has the advantage of being easier to observe in administrative data than managerial talent hoarding actions. In addition, such a test is not necessarily restricted to one specific measure of managerial hoarding behavior.
3 Setting and Data
My analysis relies on a unique combination of personnel records, internal job application data, and survey evidence from a large manufacturing firm.
3.1 Firm Overview
I collect rich data on over 30,000 white-collar and management employees from a large manufacturing firm. This anonymous firm is one of the largest manufacturers in Europe and employs over 200,000 workers around the world, the plurality of which work in Germany. I restrict my sample to Germany because it represents the largest internal labor market at the firm. The firm operates in many other countries, including the United States. The firm’s establishments outside of Germany share many features in common with those in Germany, including organizational design and internal labor market policies (e.g., application systems, widespread use of performance and potential ratings). Because I am interested in career progression to higher-level positions, my analysis focuses on employees in white-collar and management positions (i.e., employees that are either already in or could ultimately be promoted to managerial positions). There are over 200 occupations represented in this sample, ranging from technical positions in engineering to support functions in HR and finance. Table 1 describes my analysis sample, which consists of over 300,000 employee-by-quarter observations from 2015 to 2018.121212To maintain confidentiality, I do not disclose the exact number of employees in my sample. Women represent 21% of employees in the sample, stemming from the underrepresentation of women in technical occupations. Employees’ educational qualifications are high, a result of restricting the sample to white-collar and management employees. The average employee holds a Bachelor’s degree and 92% of employees work full-time. Tenures at the firm are long, with an average of 13 years, highlighting the importance of internal career progression for employees’ long-term income. Managers (i.e., those that lead a team) comprise 19% of the sample. The demographics of the employees at the firm are comparable to other large manufacturing firms in Germany. In Appendix Table B3, I compare employees in my sample to those employees in large manufacturing firms in the BiBB, a representative survey of the German workforce conducted in 2018. I find very similar patterns with respect to most employee characteristics (e.g., gender, age, German citizenship, and marital status). The firm also resembles other large firms with respect to the design of its internal labor market (hkp, 2021). As in most large German firms, employees who want to switch to a new position or to be promoted are required to apply for the position using a centralized online job portal at the firm. All job openings are posted to the job portal, where openings are visible to all employees. Applications through the portal typically take less than five minutes to complete. While employees can choose to apply to multiple positions at the same time, the median applicant applies to only one position in a given quarter. Callback and hiring decisions are also recorded in the job portal. The internal labor market is comprised of competitive openings for new positions, much like the external labor market. Only a quarter of applications are successful and the internal labor market is both spatially and interpersonally diffuse. The firm operates in over 50 cities in 250 establishments throughout Germany and one-third of internal applications are for positions in a different city. For the vast majority of applications, applicants have not previously worked with the hiring manager of the position they are applying for, indicating that application and hiring decisions are distinct in the internal labor market. Because most teams are small, consisting of one manager and six workers on average, and because most workers in a team work in very similar positions, workers who want to move up the job ladder or seek different types of jobs must leave their team. At the firm, 97% of applications are to positions outside of a worker’s current team. Thus, managers who encourage their workers to develop their career lose team members. While workers can independently apply to any internal job opening, managers’ are tasked with assisting employees in their career planning (e.g., through regular conversations, mentoring activities, assignments to trainings). However, these activities are neither directly incentivized nor monitored at the firm. The performance of all employees, including managers themselves, is regularly evaluated by their direct supervisor. Individual contributors are only rated based on their individual achievements, but manager performance ratings take into account the performance of the team. While performance ratings have modest impacts on bonus payments for workers at lower levels, the performance of the manager and her team has substantial impacts on manager pay. For instance, for managers who hold executive positions (e.g., department head or above), bonus payments represent a quarter of their total compensation. Therefore, the composition and performance of the team can have important effects on manager compensation.
3.2 Personnel Records and Application Data
My data combine the firm’s internal personnel records from 1998 to 2019 with the universe of application and hiring decisions from 2015 to 2019. The personnel records capture over 30,000 employees, and the application data cover over 16,000 job openings and over 200,000 external and internal applicants. In my main analysis sample, I restrict to employees active at the firm from 2015 to 2018, for whom I can observe outcomes through 2019, yielding a sample of over 300,000 employee-by-quarter observations. Appendix Section D.1 contains details on data construction. The personnel records contain a large set of employee characteristics, such as age, educational qualifications, and tenure. The records also contain detailed position characteristics, such as position titles, leadership responsibility, and the reporting distance to the CEO. I supplement this data with payroll data, capturing employees’ working hours, earnings, and bonus payments. The richness of these data allow me to account for factors unrelated to talent hoarding that may influence workers’ career progression. Unless otherwise noted, all analyses include the following set of controls: female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, function, division, location, full-time status, hours, and number of direct reports. In addition, the personnel records capture workers’ assignments to teams and managers. I use this information to characterize manager behavior and construct measures of manager quality based on past outcomes of managers’ team members (e.g., absenteeism, turnover). Because these data capture team assignments over time, I am also able to construct measures of managers’ formal ties to other units at the firm, by measuring whether they have previously worked in that unit. Finally, I collect information from the firm’s talent management system that allow me to measure manager actions. These include worker evaluations, such as performance and potential ratings, which are conducted by a worker’s direct supervisor. The data also capture manager actions with respect to assigning workers to internal trainings that likely increase visibility outside the team. When testing how managerial talent hoarding affects worker actions, I focus on worker applications as a key outcome. Because all job openings are posted to the centralized portal and all applications are required to be submitted through the portal, I am able to construct a panel dataset of employees’ application histories at the firm from 2015 to 2018. Through the job portal, I also observe the outcome of each application in terms of rejections, interview callbacks, and subsequent hiring outcomes. Separately measuring applications and hiring outcomes is important given that workers’ forgone job switches could be explained by low hiring success unrelated to talent hoarding.
3.3 Survey of Managers
In addition to the employee survey that was discussed in Section 2.1, I conduct a survey of managers at the firm that was designed to elicit the hoarding incentives managers face and the actions that may follow them. Managers in the firm were invited via e-mail by the firm’s HR department. Respondents were prompted to provide their perspectives on managers’ role in developing talent at the firm based on their own experiences as managers. The survey received responses from 62% of managers, yielding over 3,000 responses. Appendix Table D2 shows that managers who responded to the survey are similar in their demographics to managers in my analysis sample. The primary objective of the survey was to directly elicit the extent to which managers perceive misaligned incentives with respect to the development of talent in the firm. Managers were asked to respond to a series of questions including explicit statements with respect to potential talent hoarding. The survey also carries the advantage that it does not rely on worker reports, allowing me to more directly infer the magnitude of managerial talent hoarding in the firm. See Appendix Section D.3 for an abbreviated version of the survey instrument.
4 Evidence from Managers: Hoarding Incentives and Actions
The conceptual framework in Section 2.2 predicts that managers should hoard workers on their team due to the misaligned incentives they face. This section uses a survey of managers in the firm to provide direct evidence that managers perceive misaligned incentives that lead to talent hoarding. I then leverage a direct measure of managers’ hoarding actions using the firm’s HR data to show that managers who face stronger hoarding incentives are more likely to take hoarding actions.
4.1 Manager-Reported Hoarding Incentives and Actions
Responses to the survey of managers described in Section 3.3 reveal that managers face strong incentives to hoard talent instead of supporting workers in their career development outside the team. The majority (55%) of managers explicitly acknowledge that supporting workers’ career development entails a conflict of interest because more developed workers are more likely to leave the team (Figure 1). This conflict of interest appears to be driven by the fact that managerial talent development is not sufficiently rewarded by the firm: only 36% of managers perceive their impact on workers’ careers to be valued by the firm at least as much as their impact on immediate team performance. Yet managers can have important impacts on their subordinates’ careers: 96% of managers believe that their direct intervention, for instance through encouraging workers or increasing their visibility outside the team, has a large impact on their workers’ career development. Although the broad majority of managers (87%) indicate that talent development is an area where managers can create a high impact for the firm, only 40% report that a record of success with respect to their workers’ career development is important for their compensation and promotion prospects. This difference indicates that although managers could benefit the firm by supporting their workers’ career development, managers perceive a lack of incentive to do so. [^13] See Appendix Table B2 for survey quotations of managers describing this lack of incentive. A key disincentive to develop worker talent is that developing workers increases the risk of them leaving the team. When asked under which conditions managers would be more likely to support workers’ career development, the most common response among managers (68%) is if replacing employees who leave the team would be easier (Panel A of Appendix Figure A1). In addition, 42% of managers indicate that being able to develop workers within the team, instead of losing them to other teams, would make them more likely to engage in talent development. These findings provide direct evidence that the risk of losing team members deters managers from supporting their workers’ career development. Managers also indicate that increasing the returns to talent development would make them more likely to support workers’ careers. Most importantly, 52% of managers state receiving more recognition for developing talent as key condition. Direct impacts on managers’ promotion prospects and financial incentives, which are not common features of the internal labor market I study, were only named by 22% and 21% of managers, respectively. Managers’ survey responses indicate that managers have ample scope to impact their workers’ career development. Panel B of Appendix Figure A1 shows that when asked about specific actions they can take, managers most frequently cite making workers visible outside the team (84%) and establishing an open dialogue with workers about career development (84%). Managers also identify allowing workers to participate in trainings (69%), encouraging employees to pursue different job opportunities (62%), providing guidance for career planning (58%), and notifying workers about job opportunities outside the team (32%). These actions also accord with the survey responses provided by workers (discussed in Section 2.1), who report that talent hoarding typically occurs through managers withholding career-planning support, discouraging workers from applying to internal positions, and reducing the visibility of worker ability (e.g., by suppressing public signals of worker ability or restricting access to high-visibility trainings). Workers’ survey responses also suggest that managerial talent hoarding actions are effective in reducing workers’ career opportunities. For example, workers who report talent hoarding in form of retaliation are 5.1 percentage points (56%) less likely to have received a job recommendation by their manager ($p$=0.000), and 4.6 percentage points (13%) less likely to have received job recommendations from anyone in the firm ($p$=0.000). Together, these results point to several complementary ways in which managers can effectively impact workers’ career development. How widespread is talent hoarding? I use two complementary survey questions to assess the magnitude of managerial talent hoarding at the firm. First, my preferred measure is based on the question, “How often might leaders find themselves in situations where they need to dissuade a team member from exploring opportunities in another department due to immediate team needs or performance goals?”. The phrasing of this question was specifically designed to reflect the definition of talent hoarding in the conceptual framework: managers deliberately reducing the likelihood that a worker leaves the team. I find that 75% of managers respond that this situation occurs at least sometimes, indicating that the broad majority of managers explicitly recognize the existence of talent hoarding. My second measure is based on asking managers to report “the reasons that may prevent leaders from investing time and effort towards their employees’ career development.” Note that this question is more indirect because it asks about foregone investment rather than actively dissuading worker mobility. Nonetheless, I find that 45% of managers cite the risk of losing workers as a key reason for not investing in employees. Moreover, 66% of managers cite the need to prioritize short-term targets over long-term employee development, which accords with the responses to the first question and underscores the importance of team performance in managerial incentives. Collectively, these responses corroborate the finding that talent hoarding is a common reaction to prevailing managerial incentives.
4.2 Measuring Hoarding Actions in the Data
Measuring managers’ talent hoarding actions in the data is difficult because it typically occurs through interpersonal interactions. The survey evidence suggests that one common way in which managers limit workers’ opportunities to leave the team is by suppressing worker visibility. To capture high-stakes talent hoarding actions in the data, I identify managers’ systematic suppression of worker visibility based on a measure of worker visibility that I collect from the firm’s HR databases. Specifically, I measure systematic differences between potential ratings and performance ratings. Performance ratings are meant to provide task-specific feedback to workers about their past performance in their current position. Potential ratings are designed to inform the firm about a worker’s future potential for higher-level positions and thus are meant to identify workers who would be likely to perform well if they were promoted. Both ratings are conducted simultaneously by a worker’s direct supervisor and are very similar to common worker assessments, such as the nine-box grid, that are used by many organizations across the world (Cappelli and Keller, 2014). An important distinction between performance and potential ratings is the extent of their visibility outside of a worker’s team. Performance ratings are private signals to the worker and are not shared with other units in the firm. If a worker applies for a job in a different unit, that unit will not be able to access the worker’s past performance ratings. The firm’s HR department describes the privacy norms regarding performance ratings as similar to the strong taboo in Germany of talking about one’s salary. In contrast, potential ratings are public signals of worker talent and are widely circulated within the firm. The HR department regularly circulates lists of high-potential workers, making them highly visible outside of worker’s current team. Intuitively, a manager who wants to hoard talent should give workers lower (public) potential ratings relative to their (private) performance ratings. A survey of the firm’s employees suggests that this method of talent hoarding is commonplace. [^14] For instance, one worker states, “Supervisors suppress potential ratings because of fear that employees will leave their current position for a promotion.” (Appendix Table B1). For a given worker, the difference between performance and potential ratings may reflect worker-specific factors that are unrelated to talent hoarding; however, because managers have discretion in conducting these evaluations, comparing systematic differences between the ratings across workers can capture manager behavior. [^15] Illustrating the importance of manager discretion in this setting, Benson et al. (2021) find in contemporaneous work that managers in a large retailer overrate men’s potential compared to women’s, possibly as a reaction to men’s higher turnover risk. My measure of talent hoarding is defined by the difference between the actual and predicted potential ratings a manager assigns to their workers. Specifically, I take the residuals from the OLS estimation of the following regression of the potential rating given to worker $i$ in quarter $t$ by manager $m$ on their performance rating and other worker characteristics:
$\text{potential}_{imt}$ $=\beta_{1}\text{performance}_{imt}+\beta_{X}X_{it}+\beta_{t}+\varepsilon_{imt}$ (2)
In Equation 2, $X_{it}$ denotes the vector of controls described in Section 3.2, $\beta_{t}$ represents quarter fixed effects. I compute the average difference (over workers and quarters) between predicted and actual potential ratings for each manager. I classify a manager as having a high propensity to hoard talent if this difference is in the top tercile of the manager distribution, indicating that the potential ratings the manager assigns are systematically lower than what one would expect. Appendix Section E provides additional information on the construction and validity of this measure. I find that the measure is reasonably stable over time, supporting the systemic notion of talent hoarding the measure is meant to capture. The talent hoarding measure is also not simply a function of managers’ ability to assess talent, suggesting that differences between performance and potential ratings are unlikely to be a result of managers’ involuntary mistakes. In addition, the underrating of potential captured by this measure does not appear to represent managers’ correct assessment of workers’ future performance. Instead, managers’ suppression of the public signals of worker talent is consistent with their incentive to reduce visibility outside of the team. This measure allows me to directly test empirical predictions for the circumstances under which managers should engage in more hoarding (Section 4.3) as well as to document how talent hoarding depresses workers’ career development (Section 5).
4.3 Testing Empirical Predictions for Managerial Hoarding Actions
To provide evidence that managers purposefully hoard talent based on managers’ revealed actions, I use the direct measure of talent hoarding I construct to assess whether hoarding actions are more likely to occur when managers face greater hoarding incentives. Following the empirical predictions from my conceptual framework in Section 2.2, I relate my measure of managerial talent hoarding for each of the over 6,000 managers in my sample to the incentives these managers face. Specifically, I analyze whether managers are more likely to be classified as hoarding-prone based on the measure discussed in Section 4.2 if they face stronger incentives to hoard. To conduct this test, I estimate a linear model for managers’ hoarding propensity using an OLS regression on an indicator that managers face a strong incentive to hoard and a set of manager controls (i.e., gender, German citizenship, age, firm tenure, and education). Motivated by the conceptual framework, I construct three proxy measures that capture variation in managers’ hoarding incentives. First, the framework predicts that managers should engage in more talent hoarding if the gains from hoarding are larger. To proxy for these gains, I classify managers based on the share of their overall compensation that is based on team performance. Managers who I classify as having high financial incentives receive three times higher variable compensation relative to those with low financial incentives, implying that losing a talented worker is more financially costly to them. Second, the framework predicts that talent hoarding should be greater if a worker’s departure has a larger impact on team performance. I proxy for this dimension using the size of the manager’s team. Intuitively, smaller teams have fewer people who can compensate for a single worker’s departure. Managers are classified as having strong hoarding incentives if their team has four or less teammates, corresponding to the bottom quartile of team size at the firm. Third, the framework predicts more hoarding if managers incur lower costs of talent hoarding. I proxy for this dimension using variation in the visibility of worker talent. In functional areas where individual achievements are difficult to observe (e.g., corporate functions such as HR or marketing), it should be very easy for managers to hoard talent because the manager has a high degree of control over worker visibility. On the other hand, in areas such as engineering where many workers file patents, it should be more difficult for managers to hoard. [^16] High-visibility functions are defined as engineering, project management, and quality management. Low-visibility functions are defined as administration, finance, safety and health, HR, marketing, it, logistics, and purchasing. Managers who face greater hoarding incentives are substantially more likely to engage in talent hoarding. Panel A of Figure 2 plots the share of managers that are classified as being hoarding-prone based on my measure of suppressing potential ratings, separately by proxies of hoarding incentives. Managers who face high financial incentives are 31% more likely to hoard talent ($p$=0.000). When using team size to proxy for hoarding incentives, I find that managers in small teams are 38% more likely to engage in talent hoarding ($p$=0.000). Finally, managers in functions with low talent visibility are 13% more likely to hoard ($p$=0.002). These results show that across each of the three proxies for hoarding incentives, hoarding actions are much more common among managers who face stronger incentives to hoard, confirming the predictions from my conceptual framework. Higher levels of talent hoarding for managers facing stronger incentives are also visible in several complementary measures of hoarding actions, which corroborate the result based on potential ratings. First, I follow survey responses indicating that restricting access to high-visibility trainings is another common way through which managers are able to suppress worker visibility. In my sample, 2% of workers are assigned by their manager in a given quarter to participate in a high-visibility training. These are prestigious in-person trainings that are organized by a centralized and well-recognized department at the firm, thereby increasing worker visibility outside the team. For example, lists of training participants are widely circulated within the firm, and many workers list trainings on their CVs. I use information on which workers a manager assigns to these high-visibility trainings to construct a complementary measure of managers’ hoarding propensity. To do so, I follow the approach in Section 4.2 and identify hoarding-prone managers based on the mean deviation between expected and realized training assignments, replacing potential ratings with training assignments in Equation 2. Similar to the results using my primary measure of talent hoarding, I find that managers are more likely to engage in hoarding in the form of restricting access to high-visibility trainings if they face greater hoarding incentives. Panel B of Figure 2 shows that managers with high financial incentives to hoard are 53% more likely to engage in talent hoarding ($p$=0.000). With respect to team size, a similar pattern emerges with managers in small teams being 26% more likely to hoard ($p$=0.000). Managers who work in areas with low talent visibility are 10% more likely to hoard ($p$=0.021). Overall, these patterns are nearly identical to those using the measure of hoarding based on potential ratings. A second source of corroborating evidence comes from managers’ survey responses. The advantage of these survey-based measures is that they identify talent hoardings based on managers’ explicit statements. Panel A of Table 2 presents the share of managers who report that having to dissuade team members from exploring opportunities outside the team is a situation that occurs at least sometimes. Column 1 shows that managers with greater financial incentives are 11% more likely to report talent hoarding ($p$=0.017). Column 2 shows that managers who work in small teams are 12% more likely to report hoarding ($p$=0.000), while those that face lower hoarding costs because they work in an area with low talent visibility are 9% more likely to report talent hoarding ($p=$0.277, Column 3). Panel B of Table 2 documents similar patterns based on the share of managers who cite the risk of losing workers as a reason why managers refrain from supporting workers’ career development. Managers facing greater financial incentives are 19% more likely to report talent hoarding ($p$=0.070), while those in smaller teams are 17% more likely to do so ($p$=0.002). Managers in areas with low talent visibility are 32% more likely to report hoarding ($p$=0.010). Taken together, these results provide evidence using several complementary measures of manager actions that managers are more likely to engage in talent hoarding if they face greater hoarding incentives. Interpreted through the lens of my conceptual framework, these results confirm that more misaligned incentives lead to higher levels of talent hoarding.
5 Evidence from Workers: Effects of Manager Rotations
This section provides evidence that talent hoarding has meaningful impacts on worker outcomes. I use manager rotations as a source of quasi-random variation in workers’ exposure to talent hoarding. When a manager learns that they will move to a new position on a different team, they no longer have an incentive to hoard workers on their current team. For workers whose manager will soon rotate, this creates a temporary window of time during which they will not be subject to talent hoarding. I test for the impacts of talent hoarding using workers’ internal job applications, which represent high-stakes actions that are critical for workers’ careers within the firm. This quasi-experimental analysis provides evidence that complements the earlier analysis of managers, and offers the advantage of not restricting to a specific measure of hoarding behavior. To support the interpretation of rotation effects as being driven by talent hoarding, Section 5.1 tests predictions from the conceptual framework, while Section 5.2 discusses potentially confounding mechanisms. These empirical exercises support my interpretation based on talent hoarding and motivate the use of rotations to demonstrate the impacts of hoarding on talent allocation in Section 6. Manager rotations represent typical features of managers’ career progression and are not restricted to specific managers or teams. As in many other firms, managers switch positions within the firm to fulfill requirements for future promotions (Cullen and Perez-Truglia, 2023). [^17] I use the term manager rotation in line with previous research to refer to the event, in which a manager leaves their team for a different position within the firm. For instance, in order to become an executive, the firm requires managers to collect experience in different parts of the firm. In general, internal job switches are routine and encouraged by the firm. Consistent with the institutional environment, I empirically document that manager rotations appear orthogonal to pre-existing worker characteristics by comparing workers who are exposed to a manager rotation to those that are not (Panel A of Appendix Table 3). I find that both sets of workers are observationally similar with respect to their demographics and their past performance. Panel B of Table 3 finds a similar balance with respect to manager characteristics, such as demographics, performance, and pre-event team-level outcomes. Appendix Figure A4 uses an event study design to document the absence of pre-trends in key team-level outcomes, such as absenteeism, further corroborating that manager rotations are not restricted to teams facing specific circumstances. My main analysis uses 1,359 manager rotations that result from a worker’s direct supervisor leaving their team to make an internal job transition within the firm. [^18] Appendix Section E presents a robustness test using manager exits, which yields similar patterns. During the four-year study period, 20% of managers rotate at least once. Rotations do not occur on a regular schedule, and workers cannot easily predict when managers will leave the team; Appendix Figure A2 documents large variation in the time that managers spend in a position before rotating. The firm reports that managers typically learn about their new position two to three quarters before they rotate, at which point they inform their teams about the rotation, following a strict policy at the firm that requires them to announce their transitions as soon as possible. [^19] Survey responses, indicating that the news of internal applications spreads quickly by internal gossip, further corroborate the immediacy of these announcements since it is readily apparent when someone has violated this policy. Because managers know they will not be affected by workers leaving the team once they find out about their departure, manager rotations should temporarily reduce hoarding incentives, and thus increase worker applications. I illustrate the dynamic effects of manager rotations using a quarterly event study of applications around the quarter in which a manager rotates. I estimate a specification with worker and quarter fixed effects, binning event time dummy variables at $t=-8$ and $t=4$, and clustering standard errors at the worker and rotation level. Panel A of Figure 3 presents quarterly event study coefficients and demonstrates that manager rotations result in a distinct transitory increase in worker applications. Event time $t=0$ denotes the quarter in which a manager rotates. Application rates increase up to three quarters before the manager rotation takes place, which is when managers start to inform their teams about their departure. However, before $t=-3$, when managers do not know yet about their job rotation, trends in applications are flat. In the quarter of rotation, applications increase by 2.3 percentage points, almost doubling workers’ baseline application rate of 2.7%. As the manager’s replacement settles in, application rates return to baseline levels after one quarter. [^20] Note that applications taking a quarter to return to baseline is consistent with the fact that in a given quarter, only few job openings are available in the internal labor market, implying that not all workers who decide to apply once they find out about the manager rotation will immediately find a suitable job opening. Since manager rotations appear to have the largest impact on worker applications in the quarter of the rotation, the remainder of the analysis focuses on worker behavior in that quarter. I analyze the effect of manager rotations on worker applications in the same quarter by estimating a linear model for workers’ internal application choices using an OLS regression of the following form:
$\text{Applied}_{it}=\delta_{1}\text{Rotation}_{it}+\delta_{X}X_{it}+\delta_{t}+u_{it}$ (3)
$\text{Applied}_{it}$ and $\text{Rotation}_{it}$ are indicators that worker $i$ in quarter $t$ applies for an internal job opening and experiences a manager rotation, respectively. $X_{it}$ includes a broad set of worker and position controls. [^21] My baseline set of controls, described in Section 3.2, include female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, and number of direct reports. Appendix Table E1 documents that my results are robust to including different controls. $\delta_{t}$ represents quarter fixed effects. Robust standard errors are two-way clustered at the rotation and worker level. Manager rotations have large positive impacts on worker applications in the quarter of rotation. Column 1 of Table 4, Panel A presents OLS estimates for the effect of manager rotations on worker applications, based on Equation 3. When a manager rotates, applications increase by 2.3 percentage points, representing a 78% increase compared to the baseline application rate of 2.9%. The application effect is not a result of managers taking their subordinates with them to the new team or of workers replacing managers in their old position: 97% of applications at the firm are for positions outside of the worker’s current team and not to the manager’s new team. In addition, the observed impact does also not depend on the characteristics of the outgoing or incoming managers (Appendix Table B6). If manager rotations reflect the impact of talent hoarding, these findings imply that managerial talent hoarding deters a meaningful share of applicants from exploring job opportunities outside the team. The following subsection presents evidence in favor of this interpretation. Interpreted in this light, this finding accords with workers’ survey responses discussed in Section 2, in which 41% of workers indicate that they are afraid to apply to internal positions due to talent hoarding.
5.1 Does Talent Hoarding Drive Rotation Effects?
This section provides evidence that talent hoarding is a key mechanism underlying the impacts of manager rotations. The conceptual framework predicts that different types of workers face different magnitudes of talent hoarding, depending on factors like productivity and manager costs of hoarding. I test these predictions by comparing the effects of rotations on different groups of workers. Worker Quality.—The conceptual framework in Section 2.2 predicts that managers are more likely to hoard high-productivity workers. Therefore, when managers rotate and talent hoarding temporarily subsides, application effects should be larger for the high-productivity workers who experienced more talent hoarding than their low-productivity peers. To test this prediction, I compare the impacts of rotations on applications by worker quality. I construct a measure of worker quality as the predicted values from a regression of applicants’ internal hiring probabilities on applicant characteristics. The predicted value of this regression provides an index of worker quality for all workers, weighting worker characteristics by their importance for hiring prospects within the firm. [^22] This definition of worker quality has the advantage that it reflects the values that the firm places on worker qualifications. Robustness exercises in Appendix Section E find similar patterns when using alternative measures such as educational qualifications or past performance. For instance, while manager rotations increase applications of workers with a graduate degree by 2.8 percentage points, this increase is only 1.3 percentage points for workers without a graduate degree ($p$-value of difference=0.008). I compare workers in the top and bottom quartile of the index. Panel A of Figure 4 shows that high-quality workers experience a 4.4 percentage point increase in applications when a manager rotates, while applications among low-quality workers increase by only 0.8 percentage points. This statistically significant difference ($p$=0.000) is in line with the theoretical predictions. Departure Costs.—Another prediction is that managers hoard talent more when the costs of worker departures are larger. I test this prediction using two measures of departure costs. The first measure is the number of other workers in the team, since a larger team contains more workers that can compensate for a teammate’s departure. Panel B of Figure 4 presents effects of manager rotations for workers in the bottom and top quartiles of team size. By including detailed worker and position characteristics, this analysis compares workers with similar qualifications in similar positions, but in teams of different sizes. Applications increase by 3.5 percentage points among workers in small teams with four or less teammates, but only by 1.7 percentage points in large teams with 10 or more teammates ($p$-value of difference=0.030). Measuring departure costs using the average number of days required to fill a worker’s vacated position yields similar results. Panel C of Figure 4 compares effects for high-cost workers (more than 140 days to fill a vacancy) to low-cost workers (less than 112 days). Applications among workers who are hard to replace increase by 2.3 percentage points, compared to 1.2 percentage points for those who are easier to replace ($p$=0.107). Talent Hoarding Propensity.—The impacts of manager rotations are strongly correlated with direct measures of managers’ propensities to hoard talent. Panel D of Figure 4 compares the impacts of rotations between managers with a high versus low propensity to hoard talent, using my measure of talent hoarding based on the suppression of potential ratings (described in detail in Section 4). Applications increase by 3.4 percentage points when a manager with a high propensity to hoard talent rotates, but only by 1.4 percentage points under managers with a low propensity to hoard talent ($p$-value of difference=0.008). The framework also predicts that hoarding-prone managers have larger impacts on more productive workers. Figure 5 confirms this prediction, indicating that rotation effects of hoarding-prone managers are concentrated among high-quality workers. Risk of Retaliation—In survey results (Section 2.1), workers report fear of manager retaliation as a reason they refrain from applying. If workers are more likely to refrain from applying to positions that carry a greater risk of retaliation, applications to these positions should exhibit particularly large increases around rotations. I test this prediction by comparing workers’ applications to job openings that are closer in proximity to their current position (i.e., within versus outside their current division, functional area, and location). Intuitively, the risk of retaliation is larger if managers are more likely to find out about worker applications. Appendix Table B5 shows that manager rotations have much larger and statistically distinguishable effects on applications in close proximity to workers’ current position with respect to all three dimensions, confirming the prediction. [^23] Manager rotations could lead to an opposite prediction on proximity if they were to operate through the visibility channel instead of retaliation. However, while the suppression of visibility is a common form of talent hoarding, manager rotations do not immediately increase visibility since worker evaluations only occur one to two times a year. External Transitions.—While managers may frequently learn about workers’ applications to internal positions and possibly retaliate or interfere with those applications, this is not the case for applications to positions outside the firm. [^24] In Germany, it is very uncommon that potential employers contact the current firm to inquire about the candidate. Except for a few specific occupations, such as academics, the use of reference letters is also very uncommon, implying that managers have both little ability to learn about or interfere with external applications. Under talent hoarding, manager rotations should thus only impact internal applications and not external applications outside the firm. Because I do not observe external applications, I test this prediction by comparing the effect of manager rotations on internal job transitions within the firm and external job transitions out of the firm. Panel B of Figure 3 shows that manager rotations only increase worker transitions within the firm and not outside of the firm, even though both types of transitions trend identically in quarters prior to the rotation. Columns 1 and 2 of Appendix Table B4 confirm this finding. Manager rotations increase internal transitions by 1.1 percentage points but have negligible effects on external transitions (0.11 percentage points) compared to the identical baseline rate of 0.8% ($p$-value of difference=0.000).
5.2 Testing for Alternative Mechanisms
The analysis in the previous section is consistent with talent hoarding being a key driver of worker responses to manager rotations; however, it remains possible that worker responses could also reflect factors unrelated to talent hoarding. For instance, managers and workers may jump ship because of bad news about the future outlook of the team, or may depart after the completion of a major milestone. Other alternative mechanisms such as worker loyalty, match effects, and role-model effects could also produce the observed worker responses. In this section, I conduct a series of additional robustness checks to assess the potential importance of such alternative mechanisms. Bad news about the future or unpleasant working conditions.— I begin by testing whether a negative shock, such as bad news about the future outlook of the unit or unpleasant working conditions, could underlie the effect of manager rotations on worker applications. Under such a mechanism, one would expect to see effects on other team-level outcomes besides applications, such as team performance and morale. Appendix Figure A4 estimates an event study around manager rotations and rejects economically significant changes in key outcomes, including team-level absenteeism and performance-related bonus pay, suggesting that rotations are not correlated with deteriorating conditions within the team. Results from two placebo tests further contradict patterns that one would expect from such shocks. First, under a correlated shock that makes managers and workers want to leave the team, one would expect increased worker applications even for managers’ unsuccessful attempts to leave the team. However, Panel A of Appendix Figure A5 finds no effect of managers’ unsuccessful rotations on worker applications. Second, because coworkers are more similar in their career opportunities to a worker, one would also expect the correlation between coworker rotations and worker applications in the presence of such a shock to be larger than that between manager rotations and worker applications. Panel B of Appendix Figure A5 estimates the impacts of rotations of the most senior teammate and shows that these events have much smaller impacts on worker applications. In addition, if the documented increase in internal applications is driven by teammates jumping ship, one would expect to see at least some effect on external transitions. However, I do not find that manager rotations yield increases in exit from the firm (Panel B of Figure 3). Moreover, if teammates want to jump ship because of the long-term future outlook of the team, one should expect the increase in applications to persist for several quarters, particularly given that in the internal labor market, it can take several quarters for suitable vacancies to become available. Instead, the observed application responses are transitory. It is also unclear why workers in smaller teams, those that are harder to replace, and those under hoarding-prone managers would be more susceptible to such a shock, as suggested by the heterogeneity results presented in Figure 4.252525Managers’ hoarding propensities are quite stable over time and are unlikely to be driven by a correlated shock. Completion of a milestone or project.—If most teams were to work on fixed project schedules together, the completion of a milestone could be another key channel for correlated shocks. Several institutional features imply that such milestones are not likely to be relevant in this setting. First, manager rotations are not designed to coincide with increases in job vacancies or team dissolution that would indicate systematic rotation patterns due to common milestones in the firm. Second, the vast majority of positions at the firm are not project-based. Even in areas where project-based work is more common, projects are not normally staffed within a team, but often have employees from different teams in one project, meaning that it is not common that an entire team and their manager will finish a milestone at the same time. To empirically test for this channel, I exclude workers in jobs related to project management (where the bulk of project-based work occurs) from my event study analysis. This robustness test yields very similar application patterns (Panel B of Appendix Figure E1), suggesting that my results are not likely to be driven by project cycles. These findings are also consistent with the absence of effects on external transitions. If milestone completion were to drive responses to rotations, one would expect some effect on external transitions. Loyalty.—In principle, manager rotations could induce workers to apply if they feel a strong sense of loyalty toward their manager. While loyalty undoubtedly represents an important part of interpersonal relations in the workplace, I do not find evidence that this mechanism drives the particular increase in applications around manager rotations. Loyalty is typically assumed to compound over time; however, Appendix Figure A3 finds no evidence that rotation effects vary by the length of time a worker was exposed to the rotating manager. Instead, rotations increase applications even for workers who have been exposed to their manager for one quarter or less. While one could have also expected that loyalty is particularly strong if the worker was personally hired by the manager, Appendix Table B4 finds similar effects for workers who were hired by the rotating manager and those who were already in the team. Moreover, one would expect loyalty effects to also result in increases in external transitions. However, Panel B of Figure 3 shows that while internal and external transitions trend identically prior to the rotation, manager rotations do not yield an increase in worker’s external transition out of the firm. Worker-Manager Specific Match Effects.—Workers may have refrained from applying prior to a manager rotation because of match effects, which could occur because the manager provided favorable task assignments or offered better training opportunities. Several pieces of evidence suggest, however, that standard match effects cannot explain the increase in applications around a manager rotation. Similar to loyalty, match effects should increase worker applications for positions outside the team because the value of their default option falls when the manager leaves. However, I find no impact of manager rotations on external transitions. Under match effects, one would also expect rotations to be more impactful for workers who were hired by that manager, relative to workers who were already on the team when that manager arrived. This is because the manager would have had the ability to select new workers based on match quality. Instead, Appendix Table B4 documents that manager rotations have very similar effects on both groups. While one would also expect that match effects compound over time, I find that application effects of manager rotations do not depend on workers’ exposure time to their manager (Appendix Figure A3). Finally, it is difficult to rationalize why match effects would give rise to the observed heterogeneity across different workers, for instance in terms of larger rotation effects for workers under hoarding-prone managers. Evidence discussed in Section 4 suggests that match effects are likely smaller, not larger, under hoarding-prone managers, for example because they deter workers’ participation in trainings. Salience and Role-Model Effects.—Information-based mechanisms could also explain why rotations affect worker applications. For instance, managers who pursue rotations may generate information flows that make career planning more salient. However, since managers’ unsuccessful rotation attempts do not increase applications even though workers likely find out about them, such a channel does not appear to operate through the salience of career planning. Successful manager rotations, on the other hand, could increase applications by providing information about the success likelihood of applications. If this were the case, one would expect that the increase in applications would persist, rather than quickly dissipate, given that few vacancies are available each quarter. Role-model effects are often found to be particularly impactful if role models are similar in attributes to affected individuals (Porter and Serra, 2020, Meier et al., 2020), but homophily does not appear to play an important role in my setting. First, motivated by the previous literature on homophily by gender, I show that the rotation of a manager with the same gender yields a similar and statistically indistinguishable application response relative to the rotation of a manager with the opposite gender (Panel A of Appendix Table B6). Second, if observing others navigate their career is a key underlying factor of rotation effects, such a role-model effect should not be limited to managers. Since coworkers are more similar to workers than their managers with respect to their career opportunities, one would likely expect larger effects for observing coworkers rotate. However, Panel B of Appendix Figure A5 finds much smaller, rather than larger, effects of coworker rotations relative to manager rotations. These findings indicate that any role-model effects that are at play likely operate through channels that differ from those documented in prior research.
5.3 Discussion and Interpretation
The preceding results provide empirical evidence that the worker application responses to manager rotations are primarily driven by managerial talent hoarding. Although it is not possible to exhaustively rule out all potential alternative mechanisms, the observed patterns impose strong restrictions on the manner in which such mechanisms would have to affect worker behavior, which provides evidence against a quantitatively meaningful role for many plausible alternative mechanisms. Nonetheless, it remains possible that the observed responses to rotations may also capture factors that are unrelated to talent hoarding. Yet, to the extent that talent hoarding is a key driver of responses to manager rotations, analyzing impacts of rotations is informative about the impacts of hoarding on workers’ careers. This interpretation is supported by the analysis focusing on managers in Section 4. The following section analyzes impacts on workers’ subsequent career outcomes.
6 Consequences of Talent Hoarding
The preceding results indicate that talent hoarding effectively deters workers from applying for job opportunities outside the team. This finding naturally begs the question of whether deterred applications would have been successful, and whether talent hoarding has negative consequences for the firm overall. To answer these questions, this section measures impacts on the realized allocation of talent. In principle, talent hoarding can have negative consequences for the firm through many different channels, such as depressing human capital investment. Motivated by prior work studying the allocation of talent (Holmstrom, 1979, Rosen, 1982, Lazear et al., 2015), I focus on providing suggestive evidence that managerial hoarding affects how talent is allocated in the firm.
6.1 IV Approach
To test whether deterred applicants would have been successful, I use manager rotation as an instrument for worker applications to estimate the success probability of marginal applicants, who would not have applied in absence of a manager rotation. Equation 4 represents the reduced-form model for the effect of manager rotation on workers’ hiring outcomes:
$\text{Hired}_{it}=\theta_{1}\text{Rotation}_{it}+\theta_{X}X_{it}+\theta_{t}+\epsilon_{it}$ (4)
In Equation 4, $X_{it}$ denotes the vector of controls described in Section 3.2, $\theta_{t}$ represents quarter fixed effects. $\text{Hired}_{it}$ is a binary indicator, which is always zero if workers do not apply. Standard errors are two-way clustered at the rotation and worker level. The IV-estimator divides the reduced-form effect of manager rotation, $\theta_{1}$, by the first-stage effect on application choice, $\delta_{1}$:
$\beta_{IV}=\frac{\theta_{1}}{\delta_{1}}$ (5)
I estimate $\beta_{IV}$ by two-stage least squares, which can be interpreted as the local average treatment effect (LATE), defined as the effect of applications on hiring outcomes for workers induced to apply by manager rotations. Because workers can only get hired if they apply and there are no always takers, LATE equals the treatment effect on the treated (TOT). To interpret $\beta_{IV}$ as LATE, four assumptions are required to hold: relevance, independence, exclusion, and monotonicity (Angrist and Imbens, 1995, Angrist et al., 1996). I now provide evidence in support of these assumptions. Relevance.— Rotations have large and significant effects on worker applications (Section 5). Independence.— Manager and worker characteristics are balanced and neither worker applications nor other key team-level outcomes, such as absenteeism or bonus pay, are characterized by pre-trends, suggesting that rotations are as good as randomly assigned (Section 5). Exclusion.—The exclusion restriction requires that manager rotation does not affect hiring outcomes other than through workers’ decisions to apply. This assumption would be violated if departing managers intercede in workers’ hiring outcomes, for instance by taking workers with them to their new team or replacing themselves with workers. However, this occurs in less than 3% of applications, suggesting that departing managers generally do not make subsequent hiring decisions for their subordinates. Alternatively, departing managers may influence hiring outcomes by reaching out to hiring managers. I test for this possibility by comparing applications where the posted job is located in the same area as workers’ current job, making it more likely that the departing manager interacts with the hiring manager. However, I do not find that relative hiring rates are higher for job openings to which managers have closer ties (Panel A of Appendix Table B7). Another violation could occur if manager rotation increases worker qualifications, making them more likely to get hired. If this channel is quantitatively important, one would expect to see larger hiring effects for workers under managers with higher quality or who have been exposed to the manager for longer. I measure manager quality using past means of three outcomes: team turnover, team absenteeism, and managers’ performance rating. Panel B of Appendix Table B7 documents that the marginal hiring probability is similar for high-quality vs low-quality managers. Panel C of Appendix Table B7 shows that marginal hiring probabilities are comparable across exposure length, even for workers who have been with the manager for only one quarter, suggesting that such a channel is unlikely to violate the exclusion restriction. Monotonicity.— My results do not appear to be biased by the existence of a large defier population. In my setting, defiers are individuals who would have applied in the absence of a manager rotation, but whose unobserved propensity to apply is reduced by the instrument. Following Arnold et al. (2018) and Bhuller et al. (2020), I show that the first-stage relationship between applications and manager rotations remains positive for 16 key subgroups of workers (Panel A of Appendix Table B8). To further test for the presence of defiers, I use two proxies for workers’ unobserved application propensity: workers’ own past application behavior and the leave-out team mean of past application rates. Panel B of Appendix Table B8 shows that even workers, for whom the proxies predict that they are more likely to be defiers because of their higher application propensity in the past, do not experience lower or negative application effects when managers rotate.
6.2 Impacts on the Allocation of Talent
To document the extent to which managerial hoarding deters overall talent mobility in the firm, I analyze impacts on applications for any type of internal position. I begin by estimating whether deterred applicants would have been successful in getting hired. Column 2 of Table 4, Panel A reports 2SLS estimates of Equation 5. I find that marginal applicants face a 49.1% hiring probability. This large and positive hiring probability implies that a meaningful share of marginal applicants forgo high-stakes applications that would have been successful, indicating that the negative consequences of talent hoarding are non-negligible. This hiring probability of marginal applicants is also much higher than the average hiring likelihood at the firm, which is 27.6%, suggesting that marginal applicants are positively selected based on the assessments of the firm’s hiring managers.
A complier analysis based on Abadie (2003) corroborates that talent hoarding reduces applicant quality. [^26] Under standard IV assumptions, complier characteristics can be estimated as $E[X_{it}|\text{Compliers}]$ for some characteristic $X_{it}$. I calculate average complier characteristics and standard errors by performing 2SLS using the first-stage Equation 3 and a reduced-form equation replacing the outcome variable in Equation 4 with $X_{it}A_{it}$, where $X_{it}$ corresponds to a characteristic of individual $i$ and $A_{it}$ is a binary indicator for $i$ applying in quarter $t$. I compute characteristics for always takers, who apply even in the presence of talent hoarding, by estimating an OLS regression of $X_{it}A_{it}(1-Z_{it})$ on $A_{it}(1-Z_{it})$, which allows me to estimate $E[X_{it}|\text{Always takers}]$. Panel A of Table 6 compares average characteristics across the entire employee population (Column 1), always takers who apply even in absence of manager rotations (Column 2), and marginal applicants who only apply if a manager rotates (Column 3). I find that marginal applicants are positively selected. While 48.9% of always takers hold a graduate degree, this is true for 59.4% of marginal applicants. Similarly, 61.8% of always takers received high performance ratings prior to applying, compared to 70.4% of marginal applicants. These results are consistent with the predictions of the conceptual framework, which indicate that talent hoarding reduces the share of high-quality workers in the applicant pool, potentially limiting the ability of the firm to fill positions with suitable candidates. The fact that marginal applicants are of higher quality and that the firm would have liked to hire a high share of them suggests that these applicants would have been good fits for the new positions. To provide suggestive evidence that these applicants would have likely been productive in their new positions, I use data on employees’ performance ratings, which are designed to provide task-specific feedback on whether a worker has accomplished her tasks in the past evaluation cycle. Since most workers in a team perform very similar tasks, performance ratings are particularly well-suited for drawing comparisons across workers within teams. I assess the possibility of forgone performance by estimating a 2SLS regression in which the outcome is defined as a worker landing a position and performing better than the average performance of other workers in the new team one year later. Column 3 of Table 4, Panel A presents the results of this analysis and shows that 17.7% of marginal applicants who apply to and land a new position also outperform their teammates.
Because of the important role that allocating workers to higher-level positions plays for the efficient allocation of talent in organizations, I also analyze applications to higher-level positions, using pay as a simple measure to identify these positions. [^27] Higher-level positions are defined as those yielding a pay increase relative to workers’ current position. Robustness tests in Appendix Section E find similar results using an alternative measure to identify higher-level positions. Panel B of Table 4 shows that 20.1% of marginal applicants, who would not have applied in absence of a manager rotation, land a higher-level position and outperform their team, indicating that a non-negligible share of deterred applicants could have been productive in higher-level positions. This finding suggests that in the presence of talent hoarding, firms may be failing to fully realize productivity gains because workers are not assigned to positions in which they would be productive.
In addition to creating efficiency costs, talent hoarding may impact between-group inequality if some workers are more affected than others. The survey evidence in Section 2.1 indicates that talent hoarding occurs through direct interpersonal interactions, suggesting that impacts may be larger for workers who depend more heavily on managerial support or are more sensitive to confrontation with their manager. In line with previous work (Bertrand, 2011), employees’ survey responses highlight the possibility that talent hoarding may affect women more than men (Figure 6). Women are 26% more likely to mention the importance of manager support for their career development ($p$=0.000) and are 22% more likely to rank a good relationship with their supervisor as the most important feature of their job ($p$=0.001). These differences suggest that women rely more on managers’ career guidance and place more value on preserving a good relationship with their manager. I conduct three exercises to assess the extent of gender differences. First, I conduct the complier analysis separately by gender in order to analyze whether talent hoarding affects women at a higher part of the quality distribution. Panels B and C of Table 6 compare average characteristics across the entire employee population (Columns 4 and Column 7), always takers who apply even in absence of manager rotations (Columns 5 and 8), and marginal applicants who only apply if a manager rotates (Columns 6 and Column 9). For both men and women, marginal applicants are positively selected. However, the extent of this positive selection is more pronounced for women. While 61.2% of marginal female applicants hold a graduate degree, this is only true for 36.3% of always takers. For men, 59.4% of marginal applicants and 52.5% of always takers hold a graduate degree. Similarly, 77.7% of women at the margin received high performance ratings relative to 61.5% of always takers. Among men, 69.2% of marginal applicants versus 62.0% of always takers received high ratings. These patterns suggest that higher-quality women are affected by talent hoarding. Second, I compare mobility effects by gender by estimating the hiring probabilities for men and women at the margin. Columns 3 and 4 of Table 5 report the 2SLS estimates of Equation 5, which are separately estimated by gender and capture the probability of landing a higher-level position for marginal applicants. I find that both men and women experience positive and statistically significant marginal hiring probabilities of 41.2% and 62.3%, respectively. These hiring probabilities are much larger than those of always takers in my sample, which are 26.9% for men and 28.6% for women. Although the gender difference among marginal applicants is not statistically significant, the extent of positive selection within women is striking (i.e., marginal vs. average applicants), and suggests that talent hoarding may be more consequential for women’s career outcomes. A third test for disparate impacts by gender compares the average potential outcomes with respect to compliers’ earnings in the treated state (i.e., when marginal applicants apply because talent hoarding ceases) to those in the untreated state (i.e., when marginal applicants do not apply due to talent hoarding). This analysis leverages the potential outcomes framework that follows from the interpretation of $\beta_{IV}$ as the LATE for marginal applicants. The advantage of this approach is that it allows a comparison of the same set of individuals across two different potential outcomes, avoiding potential composition bias. Figure 7 presents estimates of log annual earnings in quarter $t+4$ for marginal applicants in both potential outcomes states by gender. Outcomes are reported in terms of percentiles at the firm. Both men and women experience higher earnings if they choose to apply; however, the larger gains realized by women lead to a reduction in the pay gap of 86%. These findings suggest that talent hoarding may exacerbate gender inequality in the firm, raising the possibility that talent hoarding has non-negligible consequences with respect to both efficiency and equity in the internal labor market.
7 Conclusion
This paper provides the first empirical evidence that talent hoarding is an important source of frictions in organizations. Using a unique combination of survey and administrative data from a large manufacturing firm, I show that managers face strong incentives to hoard talented workers in their team and act accordingly. My findings indicate that managerial talent hoarding is commonplace (explicitly reported by 75% of managers) and has non-negligible effects on the internal allocation of talent within the firm. While my results provide the first detailed insights on talent hoarding, additional evidence suggests that such talent hoarding behavior is endemic. In a survey of the top publicly listed companies in Germany, 83% cite talent hoarding as a crucial friction in their organization (hkp, 2021). Firm surveys in other countries, such as the United States, document that talent hoarding is not limited to German organizations (i4cp, 2016, KornFerry, 2015). Because talent hoarding arises due to misaligned incentives, a natural solution would be to more closely align the incentives of managers with those of the firm. In theory, firms could reward managers based on their subordinates’ future performance, irrespective of whether subordinates remain on the team. However, HR leaders typically state that realigning incentives through financial compensation is infeasible (hkp, 2021). Similarly, when asked under which conditions managers would be more likely to support workers’ career development, only a minority of managers cite financial incentives. Instead, 68% of managers highlight the ease of replacing lost talent as key condition. In addition, policies that increase application rates — such as implementing regular application schedules and having other organizational agents, such as the HR department, directly invite workers to apply for positions — could reduce the scope for managers to engage in talent hoarding. The survey I conduct shows that for such policies to be effective, the firm must be able to deter managers from retaliating against workers, for instance by assuring full confidentiality for applicants.
Although the potential benefits of these policies for efficient talent allocation may be large, it is possible that such policies could discourage managers from investing in their workers’ human capital. While a nuanced assessment of the costs and benefits associated with managerial talent hoarding incentives is outside the scope of this paper, survey responses suggest the existence of additional costs of talent hoarding that are separate from those relating to talent allocation. For example, surveyed workers who report being subject to talent hoarding are 30% more likely to report having searched for external jobs, indicating that talent hoarding may create unwanted turnover of high-quality workers the firm would like to retain. In addition, the survey responses indicate that managers hoard talent by deterring workers from pursuing training programs and from participating in high-profile projects, limiting skill development. Together, these findings suggest that managerial talent hoarding has broad impacts on talent in organizations.
Appendix A Appendix Figures - For Online Publication
Figure A1: Manager-Reported Incentives and Actions Regarding Talent Development
Panel A. Incentives for Managers to Support Workers’ Careers
Panel B. Effective Ways for Managers to Impact Workers’ Careers
Notes: This figure uses responses from the manager survey to depict manager-reported incentives and actions with respect to talent development. Panel A shows the share of managers who indicated that a given condition would make managers more likely to support workers’ career development. Panel B shows the share of managers who chose a given action as an effective way through which managers can impact their workers’ career development. Managers were allowed to choose multiple answers. See Appendix Section D.3 for more information on the manager survey and the exact question wording. 95%-level confidence intervals are displayed. Controls: Female, age, German citizenship, firm tenure, function. N=3,xxx.
Panel B. Effective Ways for Managers to Impact Workers’ Careers
Figure A2: Number of Manager Rotations by Managers’ Length in Position
Notes: This figure illustrates variation in the timing of manager rotations, as measured by the number of quarters a manager has been in their position at the time of rotation. The number of observations is 1,359, representing the total number of all manager rotations I use for my analysis.
Figure A3: Effect of Manager Rotations by Workers’ Exposure Length to Manager
Notes: This figure assesses heterogeneity in the effect of manager rotation on applications by workers’ length of exposure to the rotating manager. Each coefficient stems from a separate regression based on Equation (3). Worker subgroups are defined by the number of quarters a worker has worked under the manager. Baseline application rates are 0.031 ($\leq$1), 0.029 ($2$), 0.032 (3-4), 0.033 (5-6), 0.030 (7-10), and 0.024 (11+). The total number of observations are 3x,xxx ($\leq$1), 3x,xxx (2), 5x,xxx (3-4), 3x,xxx (5-6), 5x,xxx (7-10) and 1xx,xxx (11+). The coefficients are not statistically distinguishable from each other (the lowest p-value from pairwise comparisons is 0.2180). Robust standard errors are clustered at the worker and rotation level and 95%-level confidence intervals are displayed. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Figure A4: Effect of Manager Rotation on Team-Level and Manager-Level Outcomes
Panel A. Team Absenteeism
Panel B. Team Log(Bonus)
Panel C. Manager Performance
Panel D. Manager Log(Bonus)
Notes: This figure documents the absence of pre-trends in key team-level and manager-level outcomes before a manager rotation. Panel A shows team-level absenteeism rates. Panel B shows team-level log bonus payments. Panel C shows the manager’s performance rating. Panel D shows the manager’s log bonus payment. Event time is defined relative to the occurrence of a manager rotation. The specification includes team and quarter fixed effects, but no other controls. I bin event time dummy variables at $t=-8$ and $t=4$ and cluster standard errors at the team level. Robust standard errors are used and 95%-level confidence intervals are displayed. The mean outcome in Panels A (C) as of $t=-4$ is 0.09 (2.95). Outcome means for Panel B and D are unreported for confidentiality. N=1,xxx.
Figure A5: Placebo Tests for Effects of Manager Rotations
Panel A. Results by Outcome of Manager Application
Panel B. Results by Type of Rotating Teammate
Notes: This figure presents placebo tests for manager rotations. Estimates stem from event study regressions, in which the outcome is an indicator that the worker applied to any internal position in a quarter and event time is defined relative to the occurrence of a rotation event. The specification includes worker and quarter fixed effects, but no other controls. I bin event time dummy variables at $t=-8$ and $t=4$ and cluster standard errors at the worker and rotation level. 95%-level confidence intervals are displayed. The mean application rate as of $t=-4$ is 0.027. The sample includes those who have not experienced a manager rotation. Panel A compares a successful manager rotation (in blue) to a placebo event, in which a manager applied for an internal job rotation, but did not land the position and stayed in the team (in orange). Panel B compares a manager rotation (in blue) to the rotation of the most senior coworker (in orange).
Panel B. Results by Type of Rotating Teammate
Appendix B Appendix Tables - For Online Publication
Table B1: Workers’ Survey Quotations
Form of talent hoarding
Quotations by workers
Informal underrating
“[My] former boss made me look bad at the potential new superiors behind my back in order to keep me on the team.”
Formal underrating
“Supervisors suppress potential ratings because of fear that employees will leave their current position for a promotion.”
“Out of fear that workers will leave the team, supervisors tend to underrate employees.”
“My supervisor would certainly find out if I applied to another position at the firm and that would have a negative impact on my assessment.”
Suppressed development
“[Supervisors] keep employees in their positions by preventing further development, rejecting training courses, and increasing workloads to prevent capacity for new tasks and development measures.”
“Career development is not supported by direct supervisors, instead it is actively blocked with the goal of keeping people in their current positions.”
Soft pressure
“I decided not to apply internally … because the message communicated to me in the employee dialogue was that I can’t leave the team within three years of joining.”
“My boss strongly hinted that in order for my success until now to be considered I need to follow through on my project until the very end otherwise my efforts will not be fully taken into account.”
Threats
“… my supervisor communicated very openly in a workshop that if one of his employees applied and was not hired, his career in his current department would also be at an end.”
Retaliation
“The position should fit exactly so that negative effects of applying regarding the current manager are worthwhile.”
“Fear of negative reactions from the supervisor: I have seen it from many colleagues who openly stated that they were applying. These colleagues then received no further training and no more interesting projects. The supervisor had written them off. This made the months leading up to the final change of job very difficult.”
Notes: This table displays worker quotations regarding talent hoarding at the firm. The quotations are based on workers’ free-text responses to questions on internal career development at the firm.
Table B2: Managers’ Survey Quotations
Statement
Quotations by managers
Misaligned incentives
“Regarding the development of your subordinates, there often is a conflict of interests for managers, since the employee then usually leaves the team and the position will not be approved to be refilled again.”
“Managers have no interest in developing talents because they have no direct benefit from it.”
“Managers pursue their own goals and often prevent further development of workers, because they are not rewarded for developing talent.”
“Selfish managers are not willing to promote or recommend subordinates to other areas of the firm, even if that would add value to the firm.”
“[Middle] Managers can’t improve the situation themselves since they depend on the upper management. That is why they block the development of their subordinates.”
Risk of losing talent
“I observe that managers are not interested in letting good workers leave their current position or develop them, otherwise they would have to fill the current position again.”
“Managers don’t actively support workers in switching positions because the vacant position is usually not (immediately) filled again.”
“Positions are cut if an employee takes on a new job or even dies, so managers are afraid that their team will shrink and they don’t know how to do all the work with the rest of the team. Managers would be more open to talk to employees about developing them if they would know that they can fill the vacant position again.”
“Many managers are not necessarily interested in developing workers or helping them to get a better job within the firm, because they would lose a good worker. Switches to other areas at the firm are not encouraged, even if it would have been the right move for the worker.”
“If you are good at what you are doing, it is very unlikely that you will be suggested for a higher-level position, especially if that position is in a different team.”
“Supervisors have the policy to do whatever it takes to keep people in their team, even if this means ignoring the development of the team members.”
Notes: This table displays manager quotations regarding talent hoarding at the firm. The quotations are based on managers’ free-text responses to questions on internal career development at the firm.
Table B3: Comparison of Analysis Sample to Representative Survey of German Workforce
Analysis sample Large manufacturers German workforce
(1) (2) (3)
Female 0.21 0.21 0.45
German citizen 0.89 0.85 0.88
Age (years) 43.3 44.1 43.9
Married 0.61 0.59 0.54
Children 0.80 0.62 0.62
Education (years) 15.8 11.7 11.8
Firm Tenure (years) 13.3 16.8 13.2
Weekly hours 41.1 40.5 38.1
Team leadership 0.18 0.25 0.29
Observations 1xx,xxx 1,472 13,791
Notes: This table compares the employees in my sample to a representative survey of German workers (BIBB). Because the most recent version of the BIBB was conducted in 2018, I restrict my quarterly sample to 2018. Column 1 reports summary statistics for my analysis sample. Column 2 uses data from the BiBB survey and restricts to employees in manufacturing firms with at least 250 employees in West Germany. Column 3 uses data from the BiBB survey for all labor force participants in West Germany. Columns 2 and 3 are weighted to achieve representativeness of the German population. Team leadership is defined based on whether someone supervises direct reports.
Table B4: Effect of Manager Rotation by Transition Type and Incumbent Status
Job transition Internal application
Internal External Incumbent Newly hired
(1) (2) (3) (4)
Manager Rotation 0.0111 0.0011 0.0190 0.0275
(0.002) (0.001) (0.004) (0.004)
Outcome Mean 0.008 0.008 0.027 0.030
P-value of t-test 0.0000 0.1297
Observations 3xx,xxx 3xx,xxx 1xx,xxx 1xx,xxx
Notes: This table provides evidence in support of talent hoarding as underlying mechanism. Columns 1 and 2 illustrate the effect of manager rotations on workers’ job transitions within the firm (Column 1) and out of the firm (Column 2). Consistent with talent hoarding, only internal transitions are affected by rotations. Columns 3 and 4 document the application effects of manager rotations for workers who have been in the team before the manager arrived (Column 3) and workers who in the past were hired by the rotating manager (Column 4). I do not find that rotations have larger effects for workers who the manager was able to select herself, which is what one would expect under manager-worker-specific match effects. Robust standard errors are clustered at the worker and rotation level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Table B5: Effect of Manager Rotation by Destination of Application
Division Functional area Location
Same Different Same Different Same Different
(1) (2) (3) (4) (5) (6)
Manager Rotation 0.01587 0.00709 0.01437 0.00813 0.01768 0.00437
(0.002) (0.002) (0.002) (0.002) (0.002) (0.001)
Outcome Mean 0.01438 0.01434 0.01575 0.01310 0.01777 0.00963
P-value of t-test 0.0013 0.0214 0.0000
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx
Notes: This table illustrates heterogeneity in the effect of manager rotations on worker applications by the destination of applications. Under talent hoarding, workers should disproportionally hold off on applications about which managers are likely to find out (e.g., because of proximity to the hiring manager). I assess three dimension of proximity compared to worker’s current position: same versus different division (Columns 1 and 2), same versus different functional area (Columns 3 and 4), and same versus different location (Columns 5 and 6). Each coefficient stems from a separate regression based on Equation 3. Robust standard errors are clustered at the worker and rotation level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Table B6: Effect of Manager Rotation by Manager Characteristics
Panel A: Outgoing Manager
Manager gender Manager vs worker gender Manager experience Team absenteeism
Male Female Same Opposite Low High Low High
(1) (2) (3) (4) (5) (6) (7) (8)
Manager Rotation 0.02112 0.03554 0.02105 0.02900 0.02014 0.02691 0.02637 0.02098
(0.003) (0.009) (0.003) (0.006) (0.005) (0.006) (0.006) (0.003)
Outcome Mean 0.029 0.029 0.029 0.029 0.029 0.029 0.029 0.029
P-value of t-test 0.1092 0.2538 0.2894 0.3691
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx
Panel B: Incoming Manager
Manager gender Manager vs worker gender Manager tenure Functional area
Male Female Same Opposite Low High Same Different
(1) (2) (3) (4) (5) (6) (7) (8)
Manager Rotation 0.02151 0.02532 0.02138 0.02396 0.02511 0.02171 0.02025 0.02374
(0.003) (0.008) (0.003) (0.006) (0.005) (0.006) (0.003) (0.005)
Outcome Mean 0.029 0.029 0.029 0.029 0.029 0.029 0.029 0.029
P-value of t-test 0.6399 0.7248 0.6242 0.5324
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx
Notes: This table illustrates heterogeneity in the effect of manager rotations on worker applications by the characteristics of the outgoing manager (Panel A) and the incoming manager (Panel B). Each coefficient stems from a separate regression based on Equation 3, where the rotation event is restricted to managers with a given characteristic. Panel A compares application effects by the outgoing manager’s gender (Columns 1 and 2), manager vs worker gender (Columns 3 and 4), bottom vs top quartile of manager experience (Columns 5 and 6), and the past absenteeism rate of the manager’s team split at the sample median (Columns 7 and 8). Panel B compares application effects by the incoming manager’s gender (Columns 1 and 2), manager vs worker gender (Columns 3 and 4), bottom vs top quartile of manager tenure at the firm (Columns 5 and 6), and whether the incoming manager was in the same vs different functional area before (Columns 7 and 8). In both panels, I find no statistical difference in the effect between each pair-wise comparison as indicated by the p-value of the corresponding t-test. Robust standard errors are clustered at the worker and rotation level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Table B7: Testing for Instrument Exclusion
Panel A: IV Estimates on Hiring by Rotating Manager’s Ties to Destination
Rotating manager has worked in the job opening’s …
Division Functional area Location
Different Same Different Same Different Same
(1) (2) (3) (4) (5) (6)
Applied 0.44231 0.50917 0.31052 0.57442 0.59534 0.44680
(0.116) (0.070) (0.089) (0.076) (0.168) (0.062)
Outcome Mean 0.003 0.005 0.002 0.005 0.002 0.006
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx
Panel B: IV Estimates on Hiring by Rotating Manager’s Quality
Turnover Absenteeism Performance
Low High Low High Low High
(1) (2) (3) (4) (5) (6)
Applied 0.48749 0.53386 0.55033 0.53776 0.57052 0.60847
(0.068) (0.118) (0.104) (0.134) (0.128) (0.126)
Outcome Mean 0.008 0.008 0.008 0.008 0.008 0.008
Observations 3xx,xxx 3xx,xxx 3 xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx
Panel C: IV Estimates on Hiring by Rotating Manager’s Exposure to Workers
Exposure length in quarters
$\leq$1 2-3 4-6 7-8 9-10 11-14
(1) (2) (3) (4) (5) (6)
Applied 0.47322 0.31729 0.52088 0.45054 0.46512 0.44879
(0.220) (0.136) (0.190) (0.142) (0.207) (0.156)
Outcome Mean 0.009 0.008 0.009 0.009 0.008 0.008
Observations 3x,xxx 5x,xxx 5x,xxx 2x,xxx 2x,xxx 3x,xxx
Notes: This table finds no evidence for a violation of the exclusion restriction regarding manager rotation as instrument for worker applications. Each panel presents two-stages least squares estimates of applying on getting hired based on Equation 5. Panel A estimates hiring outcomes for applications to which rotating managers have varying degrees of formal ties. Each column uses a different split of applications based on whether the application is to a different area of the firm, where area is defined as division (Columns 1 and 2), functional area (Columns 3 and 4), and location (Columns 5 and 6). Panel B estimates hiring outcomes by rotating managers’ quality using past means for three outcomes: team turnover (Columns 1 and 2), team absenteeism (Columns 3 and 4), and manager’s performance rating (Columns 5 and 6). Each column uses a different manager type to define the rotation event based on whether the team-level outcome is in the bottom or top quartile of the team distribution. Panel C estimates hiring outcomes by managers’ length of exposure to workers. Each column uses a different split of workers based on the number of quarters the worker has been with the manager. Robust standard errors are clustered at the rotation and worker level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Table B8: Testing for Instrument Monotonicity
Panel A: Application Effects of Manager Rotations by Subgroup
Sample Observations Baseline application rate First-stage effect Standard error
(1) (2) (3) (4) (5)
Age $\leq$ 40 years 1xx,xxx 0.04011 0.02334 (0.005)
Age $>$ 40 years 2xx,xxx 0.02149 0.02083 (0.003)
Tenure $\leq$ 5 years 1xx,xxx 0.03572 0.02918 (0.006)
Tenure $>$ 5 years 2xx,xxx 0.02669 0.02019 (0.003)
Schooling $\leq$ 13 years 1xx,xxx 0.02100 0.01161 (0.004)
Schooling $>$ 13 years 2xx,xxx 0.03122 0.02575 (0.004)
Married 2xx,xxx 0.02694 0.02003 (0.003)
Not married 1xx,xxx 0.03351 0.02635 (0.005)
Parent 2xx,xxx 0.02794 0.02199 (0.003)
Non-parent 1xx,xxx 0.03351 0.02391 (0.006)
German citizen 3xx,xxx 0.02844 0.02289 (0.003)
Non-German citizen 3x,xxx 0.03395 0.01927 (0.010)
Team leadership 7x,xxx 0.02457 0.01699 (0.006)
No team leadership 3xx,xxx 0.03001 0.02337 (0.003)
High performance 2xx,xxx 0.02931 0.02785 (0.004)
Low performance 1xx,xxx 0.03066 0.01719 (0.004)
Panel B: Application Effects of Manager Rotations by Predicted Application Propensity
Never Applied Low-application High-application
applied before team team
(1) (2) (3) (4)
Manager Rotation 0.01792 0.04240 0.01813 0.02506
(0.002) (0.012) (0.003) (0.007)
Outcome Mean 0.019 0.113 0.022 0.062
Observations 3xx,xxx 4x,xxx 2xx,xxx 8x,xxx
Notes: This table finds no evidence for a violation of the monotonicity assumption regarding manager rotation as instrument for worker applications. Panel A presents first-stage effects of manager rotation on worker applications for several subpopulations of interest, as indicated by Column 1. Estimation is based on Equation 3 and conducted separately in each subpopulation. Column 2 contains the number of observations, Column 3 presents baseline application rates, Column 4 provides first-stage effects, and Column 5 contains robust standard errors, clustered at the rotation and worker level. I find that manager rotation has a positive and statistically significant first-stage effect (Column 4) for each subpopulation. Panel B presents first-stage effects of manager rotation on applying by workers’ predicted application propensity. I use two approaches to predict workers’ unobserved application propensity. Columns 1 and 2 split the sample by workers’ own past application activity. Columns 3 and 4 split the sample by whether teams’ leave-out application rates in the past were high or low. I find a significant and positive first-stage effect even for the subsets of workers who previously had a high propensity to apply (Columns 2 and 4). Robust standard errors are clustered at the rotation and worker level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Appendix C Theoretical Appendix - For Online Publication
In this section, I provide the formal derivations for predictions 4 and 5 referenced in Section 2.2.
[…]
$q(\alpha_{i},\beta_{1})$ $>q(\alpha_{i},\beta_{2})$
$\Psi\big{(}q(\alpha_{i},\beta_{1})b-c\big{)}$ $>\Psi\big{(}q(\alpha_{i},\beta_{2})b-c\big{)}$
$\implies\text{Pr}[i\text{ applies}|\beta=\beta_{1}]$ $>\text{Pr}[i\text{ applies}|\beta=\beta_{2}]$
Prediction 5. If $\alpha_{1}<\alpha_{2}$ and $\beta_{1}<\beta_{2}$implies $\frac{\text{Pr}[i\text{ applies}|\alpha_{2},\beta_{1}]}{\text{Pr}[i\text{ applies}|\alpha_{1},\beta_{1}]}>\frac{\text{Pr}[i\text{ applies}|\alpha_{2},\beta_{2}]}{\text{Pr}[i\text{ applies}|\alpha_{1},\beta_{2}]}$
[…]
$\frac{\partial}{\partial\beta}\frac{r(\alpha_{2},\beta)}{r(\alpha_{1},\beta)}$ $=\frac{\partial}{\partial\beta}\frac{\Psi(q(\alpha_{2},\beta))}{\Psi(q(\alpha_{1},\beta))}$
$=\frac{\Psi(q(\alpha_{2},\beta))\psi(q(\alpha_{1},\beta))\frac{\partial q(\alpha_{1},\beta)}{\partial\beta}-\Psi(q(\alpha_{1},\beta))\psi(q(\alpha_{2},\beta))\frac{\partial q(\alpha_{2},\beta)}{\partial\beta}}{\big{[}\Psi(q(\alpha_{1},\beta)\big{]}^{2}}$
Omitting the denominator since $\big{[}\Psi(q(\alpha_{1},\beta)\big{]}^{2}>0$ leaves to show that
$\Psi(q(\alpha_{1},\beta))\psi(q(\alpha_{2},\beta))\frac{\partial q(\alpha_{2},\beta)}{\partial\beta}$ $<\Psi(q(\alpha_{2},\beta))\psi(q(\alpha_{1},\beta))\frac{\partial q(\alpha_{1},\beta)}{\partial\beta}$
Rearranging leads to following expression
$\underbrace{\frac{\frac{\partial q(\alpha_{2},\beta)}{\partial\beta}}{\frac{\partial q(\alpha_{1},\beta)}{\partial\beta}}}_{<0}$ $<\underbrace{\frac{\Psi(q(\alpha_{2},\beta))\psi(q(\alpha_{1},\beta))}{\Psi(q(\alpha_{1},\beta))\psi(q(\alpha_{2},\beta))}}_{>0}$
Since the left-hand side of the equation is below zero and the right-hand side of the equation is above zero, it holds that $\frac{\partial}{\partial\beta}\frac{r(\alpha_{1},\beta)}{r(\alpha_{2},\beta)}<0$ for $\alpha_{2}>\alpha_{1}.$
Appendix D Data Appendix - For Online Publication
This section provides additional information on the firm records and survey data.
D.1 Dataset Construction
This paper uses a new dataset that consists of personnel records and application data from a large manufacturing firm. To comply with the firm’s data protection regulations, the data was assembled and anonymized in a secure data environment at the firm. Data from three data sources was combined: personnel records, talent management data, and job application data. First, following variables were collected from the firm’s personnel records: employment status, date of birth, gender, marital status, family status, parental leave status, date of entry into firm, highest educational degree attained, citizenship, function, division, location, working hours, number of direct reports, reporting distance to the CEO, managerial autonomy, base pay, bonus payments, absenteeism, date of exit from firm, date of retirement, id of manager, id of team. To create a quarterly panel dataset, the data were collapsed to the employee-by-quarter level, keeping the first observation of an employee in a given quarter. Second, following variables were collected from the firm’s talent management data: performance ratings, potential ratings, nominations to talent lists, worker assignments to trainings. Third, following variables were collected from the firm’s application data: application status and date, the outcome of an application (e.g., rejection, interview, offer), id of vacancy, id of hiring manager, hierarchy level of vacancy, division of vacancy, functional area of vacancy, location of vacancy. The three data sources were linked using a unique employee id within the firm’s secure data environment.
D.2 Employee Survey
All employees in my sample were invited via e-mail by the firm’s HR department and were asked to provide their perspectives on the internal labor market at the firm. The survey received over 15,000 responses, yielding a 50% response rate. Respondents are similar to employees in my analysis sample in terms of demographics and position characteristics (Appendix Table D1). Employees described challenges regarding their internal career progression both in multiple-choice answers and in the form of free-text responses. The median response time was 13 minutes. For my analysis, I only keep respondents who took at least five minutes to complete the survey. For questions including free-text responses, I exclude the minority of respondents who did not provide any answer when presenting the share of respondents who gave a particular response. The translation of the relevant questions for this study is presented in abbreviated format: A. Please rate following six statements. “Actively applying for positions at [Company] …” {I strongly agree, I agree, Undecided, I do not agree, I totally do not agree} A.1 “… increases future promotion chances.” A.2 “… does not matter since jobs are only posted proforma.” A.3 “… would cause negative consequences by my current supervisor.” A.4 “… is seen as disloyal to my current team.” A.5 “… is appropriate once employees are unsatisfied with their job.” A.6 “… should only be done after checking in with one’s direct supervisor.” B. Which job characteristics are most important to you? Please select the two most important characteristics from the following list. {Potential for training, Potential for promotion, Pay, Flexible hours, Location, Meaningful tasks, Familiar tasks, Challenging tasks, Good relationship with colleagues, Good relationship with supervisor} C. Did someone recommend an open position to you in the past 12 months? C.1 Yes C.2 No D. You indicated that someone recommended one or more open positions to you. Who recommended these positions to you? Please select all that apply. D.1 Colleagues within your team D.2 Colleagues outside of your team D.3 Your supervisor D.4 The supervisor responsible for the open position D.5 Your HR responsible D.6 Family or friends E. At the end of this survey, we are interested in your personal opinion about current challenges and potential improvements with respect to careers at [Company]. E.1 What were the reasons why you decided in the past not to apply for internal job openings at [Company]? {free-text response} E.2 What are the main challenges that you have encountered in your career development at [Company]? {free-text response} E.3 What are some of the ways that [Company] could be helpful to you as you are planning your career? {free-text response}
Table D1: Comparison of Analysis Sample to Respondents of Employee Survey
Sample Survey
(1) (2)
Female 0.21 0.24
German 0.90 0.94
Age $<$30 years 0.10 0.12
Age 30-39 years 0.30 0.32
Age 40-49 years 0.31 0.26
Age >=>=> =50 years 0.30 0.28
Tenure <=<=< = 2 years 0.10 0.14
Tenure 3-5 years 0.12 0.18
Tenure 6-9 years 0.12 0.17
Tenure >=>=> =10 years 0.56 0.51
Engineering 0.48 0.41
Finance 0.08 0.05
Marketing and Sales 0.11 0.07
Observations 3x,xxx 1x,xxx
Notes: This table compares average characteristics of employees in my analysis sample (Column 1) to employees who responded to the employee survey (Column 2). The sample of survey respondents is restricted to only contain responses who took at least five minutes to complete the survey.
D.3 Manager Survey
In order to provide direct evidence on managers’ incentives and potential talent hoarding actions, a second survey was sent to managers in my sample. The invitation procedure was equivalent to the employee survey with the firm’s HR department inviting managers to participate in the survey to provide their perspectives on being a leader at the firm. The survey received a response rate of 62%, yielding over 3000 responses. Respondents are similar to managers in my analysis sample in terms of demographics and position characteristics (Appendix Table D2). The median response time was 14 minutes. For my analysis, I only keep respondents who took at least five minutes to complete the survey. The translation of the relevant questions for this study is presented in abbreviated format: A. Where do you think leaders can create the most impact for [Company]? Please rate the impact that leaders can create. {Very high impact, high impact, low impact, very low impact} A.1 Strategic orientation A.2 Vision A.3 Operational efficiency A.4 Talent development A.5 Culture and morale A.6 Innovation B. When it comes to leaders’ compensation and promotion prospects, how important do you think is a record of success in each of the following? Please rate the importance. {Very important, important, somewhat important, not important} B.1 Strategic orientation B.2 Vision B.3 Operational efficiency B.4 Talent development B.5 Culture and morale B.6 Innovation C. What are the most effective ways that leaders can support their employees’ career development? Please select all that apply. C.1 Training: Offer and recommend relevant trainings that can help employees enhance their skills and knowledge C.2 Open dialogue: Establish a regular dialogue about career goals C.3 Career planning: Introduce a career planning process that provides clear guidance for employees to take next steps in their career C.4 Visibility: Ensure that decision-makers outside the team are aware of talented employees C.5 Encouragement: Encourage employees to pursue new job opportunities throughout the organization C.6 Information about job opportunities: Notify employees of job vacancies in other areas of the organization D. What are the reasons that may prevent leaders from investing time and effort towards their employees’ career development? Please select all that apply. D.1 Leaders’ lack of knowledge about career development opportunities, trainings etc. D.2 Limited resources to invest in employee development D.3 Risk of losing talent, because developing employees makes them more attractive to others D.4 Need to prioritize short-term targets over long-term employee development D.5 Lack of interest on the part of employees E. Under which circumstances would leaders be more likely to support employees’ career development? Please select all that apply. E.1 If the employee would develop within the team E.2 If replacing employees who leave the team would be easier E.3 If leaders who develop talent were to receive more visibility and recognition E.4 If leaders who develop talent were more likely to be promoted E.5 If leaders who develop talent would receive financial incentives F. Please rate your agreement with the following statements. {I strongly agree, I agree, I do not agree, I totally do not agree} F.1 “A leader’s direct intervention (e.g., encouragement, increasing visibility) has a large impact on employees’ career development.” F.2 “Leaders’ impact on employees’ long-term career development is valued at least as much as their impact on team performance.” F.3 “Talent development entails a conflict of interest for leaders, because more developed employees are more likely to leave the team.” G. How often might leaders find themselves in situations where they need to dissuade a team member from exploring opportunities in another department due to immediate team needs or performance goals? G.1 Never G.2 Sometimes G.3 Often G.4 Very often
Table D2: Comparison of Managers in Analysis Sample to Respondents of Manager Survey
Sample Survey
(1) (2)
Female 0.10 0.16
German 0.92 0.89
Age $<$30 years 0.01 0.01
Age 30-39 0.17 0.22
Age 40-49 0.40 0.34
Age >=>=> =50 0.42 0.43
Tenure <=<=< = 2 0.03 0.02
Tenure 3-5 0.06 0.04
Tenure 6-9 0.11 0.16
Tenure >=>=> =10 0.75 0.78
Engineering 0.40 0.33
Finance 0.08 0.05
Marketing and Sales 0.12 0.07
Observations 7x,xxx 3,xxx
Notes: This table compares average characteristics of managers in my analysis sample (Column 1) to the managers who responded to the manager survey (Column 2). The sample of survey respondents is restricted to only contain responses who took at least five minutes to complete the survey.
Appendix E Supplementary Results for Robustness - For Online Publication
[…]
$\text{potential}_{imt}$ $=\beta_{1}\text{performance}_{imt}+\beta_{X}X_{it}+\beta_{t}+\varepsilon_{imt}$
I compute the average difference (over workers and quarters) between predicted and actual potential ratings. I compute the mean deviation at the manager level to minimize the impact of idiosyncratic factors on whether a managers assign high ratings or not. This implies that my measure of talent hoarding is defined at the manager level and assigns each manager a propensity to hoard based on all performance and potential ratings they have assigned for all workers in their team during my sample period. The median manager is observed during 75% of my sample period. For those, who are present during the entire sample period, the median number of worker ratings that is used for identifying hoarding is over 70. I also find that managers’ mean deviation is reasonably stable over time, supporting the systemic notion of talent hoarding the measure is meant to capture. When using earlier years to estimate a manager’s mean deviation, its correlation with the manager’s deviation based on later years is strongly positive (0.58). Appendix Figure E2 shows the distribution of the mean deviation between predicted and actual potential ratings. The data for this figure is at the manager level (i.e., containing one observation for each of the over 6,000 unique managers in my sample). The median difference is -0.006, while the standard deviation of the difference is 0.212. I classify a manager as having a high propensity to hoard talent if the difference between the predicted and the actual rating is in the top tercile (above a cutoff of 0.1036) of the distribution across all managers. This implies that I classify over 2,$0$ managers as having a high propensity to hoard talent based on their suppression of worker visibility. In contrast, I define a manager as having a low propensity to hoard if the difference is in the bottom tercile (below a cutoff of -0.099). I verify that my results are not sensitive to this cutoff choice. Panel A of Appendix Table E4 documents similar results when using alternative cutoffs for identifying more and less hoarding-prone managers: Columns 1 and 2 split managers based on the median difference. Columns 3 and 4 use bottom and top quartiles of the difference as alternative cutoffs. Similarly, instead of assessing the effect of different manager rotations in the joint sample, Columns 5 and 6 show similar results when splitting the sample of workers by whether their manager is in the bottom and top tercile of the hoarding measure. Several additional tests suggest that my measure indeed captures managerial hoarding actions. First, the underrating of potential that the measure captures does not appear to result from managers’ involuntary mistakes. I find no statistically detectable difference in managers’ hoarding probability based on the extent of experience they have in leading (and rating) a team. The underrating also does not appear to result from managers’ accurate assessment of worker potential. When managers with high propensities to hoard talent rotate, underrated workers not only experience increases in applications and hiring, but are also likely to perform well in subsequent positions, demonstrating that the low potential rating was inaccurate. In addition, while low potential ratings could in theory stem from the fact that managers have an incentive to hire low-potential workers to avoid the possibility of losing talent, this is not confirmed in the data. I find that talent hoarding effects occur both for incumbent workers and workers who are newly hired by a rotating manager (Columns 3 and 4 of Appendix Table B4) . A placebo test provides further evidence against the importance of alternative channels that may be underlying the systematic difference between performance and potential ratings. While managers are likely to learn about workers’ unsuccessful applications within the firm and have some power to diminish workers’ application probability, it is unlikely that managers will be able to observe or intervene with respect to applications outside of the firm. Consequently, under talent hoarding, comparing managers with high and low propensities to hoard talent should lead to differential effects with respect to workers’ internal career progression, but to zero impacts on transitions to jobs outside of the firm. However, if managers’ suppression of potential ratings is not a sign for talent hoarding, but reflects other types of manager-specific behavior that affects workers, we would not necessarily expect the effect on external transitions to be zero and the effect under both high-propensity and low-propensity talent hoarders to be similar. I use this intuition to conduct a placebo test, comparing rotation effects for managers with high versus low propensities to hoard for internal applications, internal job transitions within the firm, and external job transitions out of the firm. Appendix Table E3 documents that I find a zero effect on external transitions for both managers with high and low propensities to hoard talents (Columns 5 and 6), which contrasts my findings on internal applications (Columns 1 and 2) and internal job transitions (Columns 3 and 4). In addition to testing robustness of my measure of talent hoarding, I also verify the robustness of the measure of worker quality I use for my heterogeneity analysis. While my preferred approach uses the bottom and top quartile of the quality index I construct, Panel B of Appendix Table E4 documents similar results when using below and above the median as alternative cutoff points (Columns 1 and 2). My results are also robust to alternative measures of distinguishing between high-quality and low-quality workers. Columns 3 and 4 split the sample of workers by the median years of education they received. Columns 5 and 6 split the sample by the median performance rating. The resulting application patterns are very similar to my preferred measure of worker quality.
Figure E1: Effects of Alternative Rotation Events
Panel A.Manager Exits as Event
Panel B. Sample Without Project Management
Notes: This figure presents robustness tests for the mechanisms underlying manager rotations. Panel A focuses on manager rotations that were induced by the outgoing manager exiting the firm. Panel B focuses on manager rotations based on internal switches, but excludes workers from the analysis sample who work in project management positions and are most likely to work in projects with fixed milestones. Estimates stem from event study regressions, in which the outcome is an indicator that the worker applied in a quarter and event time is defined relative to the occurrence of a rotation event. The specification includes worker and quarter fixed effects, but no other controls. I bin event time dummy variables at $t=-8$ and $t=4$ and cluster standard errors at the worker and rotation level. 95%-level confidence intervals are displayed. The mean application rate as of $t=-4$ is 0.027. The sample includes those who have not experienced a manager rotation.
Panel B. Sample Without Project Management
Table E1: Robustness in Application Effects
No Employee Position Baseline Baseline +
Controls Demographics Controls Controls Rating + Mobility
(1) (2) (3) (4) (5)
Manager Rotation 0.0222 0.0220 0.0228 0.0226 0.0234
(0.003) (0.003) (0.003) (0.003) (0.003)
Outcome Mean 0.0290 0.0290 0.0290 0.0290 0.0290
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx
Notes: This table presents results from robustness exercises using different sets of controls to estimate the effect of manager rotations on worker applications. Column 1 presents the raw estimates, which only control for quarter fixed effects. Column 2 controls for quarter fixed effects and employee demographics, including gender, age, German citizenship, educational qualifications, marital status, and family status. Column 3 controls for quarter fixed effects and for employees’ position characteristics (i.e., function, division, location). Column 4 contains my baseline set of controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects. Column 5 adds controls for worker evaluations (i.e., performance and potential ratings) and workers’ past internal mobility. Robust standard errors are clustered at the worker and rotation level.
Table E2: Robustness in Effects of Manager Rotation on Worker Outcomes
Applied Hired Hired + Perform
OLS IV IV
(1) (2) (3)
Manager Rotation 0.0113 - -
(0.002) - -
Applied - 0.4675 0.1782
- (0.079) (0.060)
dd
Outcome Mean 0.0118 0.0027 0.0013
Observations 3xx,xxx 3xx,xxx 3xx,xxx
Notes: This table reports the effects of manager rotations on workers’ career outcomes with respect to applying and landing higher-level positions. In this table, higher-level positions are defined as those that have a hierarchy ranking of at least 10 based on methods developed in Haegele (2024). Each coefficient is based on a separate regression. Column 1 reports the first-stage effect of manager rotation on applications based on Equation 3. Column 2 reports estimates from a two-stages least squares regression on landing a position that instruments for applying with manager rotation based on Equation 5 which represents the LATE. Column 3 estimates a similar two-stages least squares regression, but use an indicator for landing a position and performing better than the leave-out team average one year later as outcome variable. Robust standard errors are clustered at the worker and rotation level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Figure E2: Mean Deviation Between Actual and Predicted Potential Ratings
Notes: This figure depicts the mean deviation by manager between predicted and actual potential ratings, which captures systematic discrepancies in worker visibility, and serves as a measure of managers’ propensities to hoard talent. The y-axis reports the number of unique managers with a given deviation. I assess managers’ systematic underreporting of public potential ratings by comparing managers’ actual potential rating to the predicted potential rating based on managers’ own assessment of worker performance and worker characteristics. Values above zero represent managers who on average lower their public signal below their private signal of worker talent, which captures one likely dimension of talent hoarding. I use the top (marked in orange) and bottom (marked in blue) tercile of this distribution to classify managers as high-propensity versus low-propensity to hoard talents. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects. N=6,xxx.
Table E3: Placebo Test for Talent Hoarding Measure
Panel A Panel B Panel C
Internal Applications Internal Transition External Transitions
High Low High Low High Low
(1) (2) (3) (4) (5) (6)
Manager Rotation 0.0344 0.0144 0.0166 0.0101 -0.0014 0.0020
(0.006) (0.005) (0.004) (0.003) (0.002) (0.002)
Outcome Mean 0.029 0.029 0.008 0.008 0.008 0.008
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx
Notes: This table provides a placebo test for my measure of managers’ propensities to hoard talent based on the mean deviation between actual and predicted potential ratings. Each coefficient stems from a separate regression based on Equation 3. Each panel compares rotations of managers with high versus low propensity to hoard. The outcomes of interest are internal applications (Panel A), internal job transitions within the firm (Panel B), and external job transitions out of the firm (Panel C). Robust standard errors are clustered at the worker and rotation level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Table E4: Robustness in Hoarding and Quality Measure
Panel A: Measure of Managerial Talent Hoarding Propensity
Joint Sample Joint Sample Separate Sample
Bottom $\frac{1}{2}$ Top $\frac{1}{2}$ Bottom $\frac{1}{4}$ Top $\frac{1}{4}$ Bottom $\frac{1}{3}$ Top $\frac{1}{3}$
(1) (2) (3) (4) (5) (6)
Manager Rotation 0.0167 0.0287 0.0155 0.0320 0.0189 0.0323
(0.004) (0.004) (0.004) (0.005) (0.005) (0.006)
P-value of t-test 0.0771 0.0251 0.0084
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 1xx,xxx 1xx,xxx
Panel B: Measure of Worker Quality
Quality Index Education Performance
Bottom $\frac{1}{2}$ Top $\frac{1}{2}$ Bottom $\frac{1}{2}$ Top $\frac{1}{2}$ Bottom $\frac{1}{2}$ Top $\frac{1}{2}$
(1) (2) (3) (4) (5) (6)
Manager Rotation 0.0115 0.0349 0.0128 0.0277 0.0172 0.0278
(0.003) (0.005) (0.004) (0.004) (0.004) (0.004)
P-value of t-test 0.0000 0.0082 0.0628
Observations 1xx,xxx 1xx,xxx 1xx,xxx 1xx,xxx 1xx,xxx 2xx,xxx
Notes: This table documents the robustness of the heterogeneity analysis with respect to the application effects of manager rotations. Each coefficient stems from a separate regression based on Equation 3. Robust standard errors are clustered at the worker and rotation level. Panel A documents the robustness of the measure of talent hoarding based on deviations between managers’ actual and predicted potential ratings. Columns 1 and 2 split managers based on whether their hoarding measure is below or above the median instead of terciles used in my preferred measure. Columns 3 and 4 split managers based on the bottom and top quartiles of the hoarding measure. Columns 5 and 6 use terciles, but instead of classifying rotation events based on managers’ hoarding measure, I split workers into separate samples based on their managers’ hoarding measure. Panel B documents the robustness of my measure of worker quality. Columns 1 and 2 split workers into high-quality versus low-quality using the median of the quality index, instead of using quartiles as in my preferred approach. Columns 3 and 4 distinguish high-quality versus low-quality workers based on the median education level (having 18 or more years of schooling versus less than 18 years of schooling). Columns 5 and 6 make the quality distinction based on the median performance rating. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.